A Complete Bibliography of Publications in the British Journal for the History of Science: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

17 May 2024 Version 1.16

Title word cross-reference

#BSHSGlobalHist [HLDS20].

$103.00 [Amu23]. $112.50 [Tar24]. $120.00 [Fig21]. $125.00 [Bro20]. $150.00 [Dun23]. $19.00 [Win23a]. $19.95 [Dua21, Ste22]. $210.00 [Smi23]. $24.95/£ [Hel20]. $25.00 [Sim22]. $27.50 [HA21, Sac21]. $28.95 [BR23, vD23]. $29.95 [Dua21, Goo20, Ste22]. $30.00 [Hol20, Kat20]. $32.00 [Hay23]. $32.95 [Bla20]. $34.95 [Ris20]. $35.00 [Bud24, HA20, Jon21, LZ24, Mor21a, Pal21, Sho22, Smi21b]. $39.95 [Bim20]. $40.00 [Amb21b, Fig21, Hol21, McC23]. $44.95 [Tal21]. $45.00 [Lan22, Leh23, Pal21, Win23b]. $455.00 [Hor23]. $49.95 [Fra20, Hem22, Web23]. $50.00 [Amb21a, Far21, Han20, Hes23, Sle20, Spi23]. $52.00 [vD23]. $54.95 [GS21]. $55.00 [Bru24, Buc21, Hab24, Lor24]. $57.00 [Ols23]. $60.00 [Boo23a, Mül24, Sec22, Tal21]. $64.95 [Web20]. $65.00 [Bru22, Gro24, Mox21]. $79.95 [Led21]. $82.50 [HA21]. $90.00 [Mac20]. $99.00 [Chu23].

. [Mil21, Sle20, Web20].

10.14324 [Ran23]. 10.14324/111.9781800082328 [Ran23]. 111.9781800082328 [Ran23]. 1500 [Ols23]. 1660 [Exl23]. 1700 [Dun22]. 1715 [Buc21]. 1720 [Fig21]. 1820s [Ame20]. 1830s [Ame20]. 1840 [Far20]. 1856-February [Bro20]. 1859 [Bro20]. 1880 [Bla20, Web20]. 1880-1950 [Wil21a]. 1890 [Dol21]. 1900 [Hab24]. 1910 [Sle20]. 1929 [For23]. 1948 [Sha22]. 1950 [Smi21b]. 1955 [Bru24]. 1960s [Pie20]. 1970s [Pie20, Gro24].

20.00/$39.99 [Wil20]. 2015 [Ran23]. 28.00/$35.00 [Dun22].

32.50/£ [Dee20]. 35.00/$38.00 [Tim20].

62 [Kan20]. 69.99/£ [Tar22].

978-0-00-829995-8 [Wil20]. 978-0-19-881862-5 [Kor21]. 978-0-19-882352-0 [Ros20]. 978-0-226-35542-9 [Dee20]. 978-0-226-50929-7 [Buc21]. 978-0-226-54608-7 [Hol20]. 978-0-226-55323-8 [Pal21]. 978-0-226-59226-8 [Mox21]. 978-0-226-60201-1 [HA20]. 978-0-226-60568-5 [Hay23]. 978-0-226-62210-1 [Tal21]. 978-0-226-65088-3 [Hol21]. 978-0-226-68072-9 [Fig21]. 978-0-226-68086-6 [Fig21]. 978-0-226-69304-0 [Smi21b]. 978-0-226-70615-3 [HA21]. 978-0-226-70629-0 [HA21]. 978-0-226-70680-1 [Far21]. 978-0-226-71070-9 [Dun22]. 978-0-226-73160-5 [Leh23]. 978-0-226-73188-9 [Sac21]. 978-0-226-73286-2 [Sim22]. 978-0-226-74129-1 [Bru22]. 978-0-226-75250-1 [Pal21]. 978-0-226-76138-1 [Lan22]. 978-0-226-82228-0 [Tar24]. 978-0-226-82326-3 [Bru24]. 978-0-226-82328-7 [Hab24]. 978-0-2260-2038-9 [Sho22]. 978-0-262-04514-8 [Law21a, Law23a]. 978-0-262-04553-7 [Mül24].

978-0-262-53905-0 [Amb21a]. 978-0-262-53914-2 [Jon21]. 978-0-262-54361-3 [Gro24]. 978-0-2620-3961-1 [Ste22]. 978-0-2620-4514-8 [Law23b]. 978-0-2625-3897-8 [Ste22]. 978-0-300-22383-5 [Mor21a]. 978-0-300-22741-3 [Amb21b]. 978-0-300-23006-2 [Mid21]. 978-0-300-23046-8 [Kat20]. 978-0-300-23610-1 [Hol22]. 978-0-300-24373-4 [Mil21]. 978-0-300-24677-3 [Jam21]. 978-0-367-22842-2 [Jac21]. 978-0-5000-5200-6 [Bow22]. 978-0-511-98014-5 [Sia21]. 978-0-674-28136-3 [Gro23]. 978-0-674-91919-8 [Far23]. 978-0-674-97201-8 [Har21]. 978-0-6745-0462-2 [BR23]. 978-0-6747-3758-7 [Bow22]. 978-0-691-17507-2 [Des20]. 978-0-691-18188-2 [Goo20]. 978-0-7735-5312-5 [Hem22]. 978-0-8173-1985-4 [Tal21]. 978-0-8229-4511-6 [Bim20]. 978-0-8229-4533-8 [Bro20]. 978-0-8229-4551-2 [Han20]. 978-0-8229-4552-9 [Sle20]. 978-0-8229-4576-5 [Win23b]. 978-0-8229-4579-6 [Hes23]. 978-0-8229-4609-0 [LZ24]. 978-0-8229-4649-6 [Spi23]. 978-0-8229-4659-5 [McC23]. 978-0-8229-4681-6 [Sec22]. 978-0-8229-4706-6 [Bud24]. 978-0-8229-4734-9 [Lor24]. 978-0-8229-4735-6 [Hor23]. 978-0-8229-4737-0 [Boo23a]. 978-0-8229-6563-3 [Ris20]. 978-0-9955462-9-5 [Ree21]. 978-1-009-12311-2 [Arn23]. 978-1-032-13916-6 [Ant24]. 978-1-032-16223-2 [Jor23].

978-1-032-33718-0 [Amb24]. 978-1-032-53264-6 [Smi24]. 978-1-107-08872-6 [Bow23b]. 978-1-108-34823-2 [Tar22]. 978-1-108-41828-7 [Far20]. 978-1-108-41869-0 [Smi23]. 978-1-108-47819-9 [Tur21]. 978-1-108-49949-1 [Sha22]. 978-1-108-68845-1 [LD21]. 978-1-108-72884-3 [Coz20]. 978-1-108-74709-7 [Tur21]. 978-1-108-78126-8 [Dol21]. 978-1-108-84483-3 [Amu23]. 978-1-108-87915-6 [Ros22]. 978-1-1084-8713-9 [Di 23]. 978-1-1087-3219-2 [Di 23]. 978-1-1088-4457-4 [Kri23]. 978-1-1385-7015-3 [Dur23]. 978-1-316-88243-6 [Hun23]. 978-1-4214-2531-3 [Led21]. 978-1-4214-2552-8 [Fra20]. 978-1-4214-2673-0 [Dua21]. 978-1-4214-2889-5 [Web20]. 978-1-4214-2951-9 [vD23]. 978-1-4214-3852-8 [Ols23]. 978-1-4214-3908-2 [Dua21]. 978-1-4214-4527-4 [vD23]. 978-1-4426-3017-8 [Mac20]. 978-1-4696-4844-6 [Bla20]. 978-1-4696-5609-0 [Chu23]. 978-1-5017-4015-2 [Web23]. 978-1-5261-3138-6 [Kan20]. 978-1-78023-997-2 [Mey20]. 978-1-78735-393-0 [Bee21]. 978-1-78735-394-7 [Bee21]. 978-1-78735-395-4 [Bee21]. 978-1-7891-4455-0 [Hol23a]. 978-1-78914-008-8 [Hel20]. 978-1-78914-214-3 [Tur21]. 978-1-78914-299-0 [Des21]. 978-1-78914-314-0 [Eic21]. 978-1-78914-342-3 [Smi20]. 978-1-78914-440-6 [Win23a]. 978-1-78962-081-8 [Guf21].

978-1-78962-082-5 [Guf21]. 978-1-8000-8234-2 [Ran23]. 978-1-80008-415-5 [Koh23]. 978-1-85124-553-6 [Mar20]. 978-1-935408-36-9 [Hay21b]. 978-1-9788-2021-0 [Dun23]. 978-2-0211-7599-8 [Bow23a]. 978-2-503-58806-3 [Nia24]. 978-3-030-78972-5 [Exl23]. 978-3-8353-3680-3 [Sto23]. 978-94-6301-434-2 [Hol23b]. 978-9500845-399-0 [Man24]. 978-981-322-868-9 [McM20]. 978-981-327-903-2 [Tim20]. 978-987-628-598-8 [Man24]. 987-1-1088-3708-8 [GS22]. 987-1-4214-3874-0 [GS21].

Academy [Nia24]. access

[Bee21, Koh23, Ran23]. accurate [Sto22]. Achille [Sta22]. Adaptation [Law21b]. Addiction [Sac21]. Address [Boo23b]. Adjusting [Low21]. Adler [Har21]. aerodynamic [Spi20]. Aesthetic [Fig21]. Aesthetics [Amb24]. afford [Hua21]. Africa [GS22]. African [dB20]. After [vD23]. against [Med23]. Age [Des20, Des21, Hay21a, Sec22, Ker20, Sim22]. Ahmed [Tek22]. Aileen [Ran23]. Aires [Man24]. airliner [Spi20]. Airy [Ish20]. Alabama [Tal21]. Alaska [Sch22a]. Alchemy [Dun22]. alcohol [ASW21]. Alex [Pal21, Sho22]. Alexander [Fig21, Sia21]. Alexis [Exl23]. Alfred [Tal21]. Alfredo [Hol23b]. Alida [Ols23]. Alison [Exl23]. Allan [Ste22]. allowances [Ayl21]. alternatives [BS20]. Amanda [Tur21]. Ambition [Wyn21]. ambitions [Tur23]. Ameghino [Man24]. Amelia [Han20]. America [Goo20, Sac21, Sec22]. American [Bim20, Bud24, Dun23, Fra20, Jon21, Led21, Bei21]. Amir [Sha22]. analogy [Sch22b]. ancêtres [Bow23a]. Ancient [Sia21, Spi23, KW21]. Anctil [Hem22].

Andreas [Boo23a]. Andrei [Hay21b]. Andrew [Kor21, Dun23, Jor23, Mil21]. Andrews [Smi20]. Animal [Dee20, Hol20, Hol23c]. Animals [Tar24]. Anna [Nia24]. Anning [Ree21]. anthrax [Sta22]. Anthropocene [Hay21a, LZ23, Sch22a]. Anthropozoic [LZ23]. Anti [Ros23]. Anti-voluntarism [Ros23]. Anticipating [Car21]. Antiquity [GS21]. Antonio [LZ23]. Antony [Har21]. Anxiety [Dun23]. Anxious [Han20]. Appetite [Smi21b]. appliances [KW21]. appreciation [Smi21a]. approach [Wid23]. AR$3 [Man24]. AR$9 [Man24]. Arabic [KK22]. arbres [Bow23a]. Archaeology [Ker20]. Archer [Guf21]. Archibald [Win23a]. Architecture [Hor23]. archives [MS21]. Arctic [McC23]. Argentina [Amu23, dA20]. Argentinos [Man24]. Arias [Mox21]. Aristotle [Jor23]. Armed [Led22]. Aronova [Lan22]. Art [Hay21b, Smi24, BRSW23, EW23, Mau22]. Artisanal [Amb21b]. Arts [Amb21b, Jon23]. Asia [Car21]. Asian [Ano21h, Wil21a]. Assembling [Bow22]. Assessment [HA20]. Astrologers [Pfe21]. astrology [Pfe21]. Astronomer [Mül24]. astronomical [Gil20]. Astronomy [Sle20, Tur23]. Astrostereon [Bie20]. Athenæum [Jam21]. Atlantic [Far20, Ols23]. atmosphere [Fra21]. atmospheres [Ano21h, Wil21a]. atmospheric [Ame20]. atom [Ada23]. atomic [Ker20]. attempt [Ame20]. auctions [Mac22]. Audra [Dua21]. Australian [Win23b]. authority [Dis24, Kir20]. Authorship [Pal21]. automatic [Ish20]. Avant [Smi24]. Avant-Garde [Smi24]. aviation [Spi20].

B [Buc21, Mil21, Ris20]. Back [Ano20a, Ano20c, Ano20e, Ano21b, Ano21d, Ano21f, Ano22a, Ano22c, Ano22e, Ano22g, Ano23a, Ano23c, Ano23e]. background [Led22]. balance [Hol23c]. Baltimore [Dua21, Fra20, GS21, Led21, Ols23, Web20, vD23]. Banks [Mor21a]. Baranski [Lor24]. Barnett [vD23]. Barrett [Kri23]. Barton [Bro20]. Battles [Amb21a]. Bay [Mor21a]. Beatriz [Man24]. Beauty [Amb24]. Beech [Tim20]. Before [Buc21]. Behavior [Hol20]. Beijing [Tim20]. Beinart [GS22]. Being [Bee21]. Belknap [Far23]. Belyakov [McM20]. Ben [Chu23, Far20]. Benito [Mox21]. Benjamin [Gor22, Wil20]. Bennet [Jac21]. Benson [HA21]. Benton [Bow22]. Berlin [Kai22, Mac22]. Bernard [Jac21]. Bertucci [Amb21b]. between [Hol23b, Kem23]. Bible [Bie20]. Biblical [Mox21]. Biddell [Ish20]. Big [Sac21]. biologists [Lon20]. Biology [Gro23, Hop22, Hor23]. biomedicine [McG21]. birds [Nix24]. Birmingham [Wyn21]. birth [Hop22, Kom24, Wes21]. BJH [Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21a, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e]. Black [Gro23]. Blankaart [Kle20]. Block [Boo23a]. Blood [Jor23]. bodies [ASW21]. Bodleian [Mar20]. Bolivia [Chu23]. Bolshevik [Lan22]. Bonea [Han20]. Book [Amb21a, Amb21b, Amb24, Amu23, Ant24, Arn23, Bee21, BR23, Bim20, Bla20, Boo23a, Bow22, Bow23a, Bow23b, Bro20, Bru22, Bru24, Buc21, Bud24, Chu23, Coz20, Dee20, Des20, Des21, Di 23, Dol21, Dua21, Dun22, Dun23, Dur23, Eic21, Exl23, Far20, Far21, Far23, Fig21, Fra20, GS21, GS22, Goo20, Gro23, Gro24, Guf21, Hab24, Han20, Har21, Hay21b, Hay23, Hel20, Hem22, Hes23, Hol23a, Hol20, Hol21, Hol22, Hol23b, HA20, HA21, Hor23, Hun23, Jac21, Jam21, Jon21, Jor23, Kan20, Kat20, Koh23, Kor21, Kri23, Lan22, Law21a, Law23a, Law23b, Led21, Leh23, LD21, Lor24, LZ24, Mac20, Man24, Mar20, McC23, McM20, Mey20, Mid21, Mil21, Mor21a, Mox21, Mül24, Nia24, Ols23, Pal21, Ran23, Ree21, Ris20, Ros22, Ros20]. Book [Sac21, Sec22, Sha22, Sho22, Sia21, Sim22, Sle20, Smi20, Smi21b, Smi23, Smi24, Spi23, Ste22, Sto23, Tal21, Tar22, Tar24, Tim20, Tur21, Web20, Web23, Wil20, Win23a, Win23b, vD23]. Books [Des21, Eic21, Hay21b, Hel20, Hol23a, Mey20, Smi20, Tur21, Win23a, EW23]. Borch [Hol23a]. Borch-Jacobsen [Hol23a]. Border [Amu23]. Boris [Wor22]. botanical [Kai22]. Botany [Mar20, Mor21a]. Bowl [Wid23]. Box [Gro23]. Boy [Wil20]. Boyle [Leh23]. Brain [Law21b]. brains [Boo23b]. Bravo [Hel20]. Brazil [Amu23]. Bread [Mid21]. Breeding [Tar24]. Brepols [Nia24]. Breyfogle [Ris20]. brig [Bek20]. Britain [Far21, Ayl21, Han20, Hol22, Hun23, Jon23, Lon20, McA22, Sta22]. British [Kan20, Ros22, Sec22, Dav21, DUS23, Hua21, Spi20, Ste24, Wyn21]. broadcasting [Pie20]. Broughton [Mac20]. Brown [Coz20]. Browne [Bro20]. Bruce [Led21]. Brunswick [Dun23]. Brysse [HA20]. BSHS [BS20]. Bud [Bee21]. Buenos [Man24]. building [Bur22]. Bulletin [Ada23]. Burkhardt [Bow23b]. Burnet [Ros23].

C [Gro24, Hol20, Ols23, Ste22, Ano21h, Bec20, Dol21, Ish20, Pfe21, Wil21a, Wil21b, Wor22, tH23]. C. [Led21]. Cain [Goo20]. calculation [Sto22]. Cambridge [Amb21a, Amu23, Arn23, BR23, Bow22, Bow23b, Coz20, Di 23, Dol21, Far20, Far23, GS22, Gro23, Gro24, Hun23, Kri23, LD21, Mül24, Ros22, Sha22, Sia21, Smi23, Tar22, Tur21, Har21, Jon21, Pal21, Ste22]. Camilla [Ran23]. campaign [Ayl21]. can [Hua21]. cancer [Med23]. Cañizares [Ant24]. Capitalism [Mil21]. capitalist [Gil20]. Caribbean [Kan20]. Caribou [Sch22a]. Carolina [Bla20, Chu23]. Caroline [Guf21].

Cartography [Hay23]. Cartoon [DUS23]. case [Kir20]. Cassidy [Ste22]. Catholicism [dA20]. causes [Ame20]. cave [Mat22]. centre [Wyn21]. Century [Han20, Hor23, Jon21, Pal21, Dis24, Jon23, KK22, Kir20, Mac22, McA22, Pfe21, Spi20, Sta22, Sto21, Wyn21, Bee21, Far21, Hay21b, LZ24, McC23, Ros22]. Cesarean [Fra20]. Cevdet [Tek22]. Chair [Mül24]. Chakrabarti [GS21]. Chakrabarty [Sim22]. challenge [Gil20]. chamber [Ber21]. Change [Chu23, Hay21a, YJ21]. Chapel [Bla20, Chu23]. chapter [Tek22]. Charles [Bow23b, Wil20, Win23a, Tal21]. Charlotte [Tur21]. chemistry [Käl22, Som23, Hol23b]. Chennai [Tim20]. Chernobyl [McM20]. Chicago [Bru22, Bru24, Buc21, Dee20, Dun22, Far21, Fig21, Hab24, Hay23, Hol20, Hol21, HA20, HA21, Lan22, Leh23, Mox21, Sac21, Sho22, Sim22, Smi21b, Tal21, Tar24]. China [Mil21, Kri23]. Chinese [Bar23, Fra21]. Chris [Mar20]. ciné [Lon20]. ciné-biologists [Lon20]. cinema [tH23]. circa [Ols23]. Circle [Ros20]. Circuit [Sec22, Wid23]. Circulation [Jor23, Sto21]. Claire [Exl23, Hol22]. Clarke [Kan20]. clash [Zai23]. Classroom [Dun23]. Claudine [Bow23a]. clearing [Kai22]. Clements [Bim20]. Climate [Hay21a, Sim22, ASW21, RBB^{+}20]. clinical [Ste24]. cloth [Bru24, Dun23, Hab24, Leh23, Tar24, Win23b]. cloud [Ber21]. Club [Bei21, Jam21]. co [Lon20]. co-production [Lon20]. Coercion [Far21]. Cohen [Bow23a]. Cold [Dua21, Goo20, Kor21, Kri23, Lan22]. Colder [Hol21]. Collard [Tal21]. Collecting [Dol21, Nia24, McA22]. collection [Mac22]. Collective [Nia24]. Collins [Wil20]. colonial [Kai22, Web23, Wil21b]. Colouring [EW23]. Commerce [Far21]. Commercial [Kem23]. committee [Käl22]. Commodification [Hes23]. Common [Hes23]. Communicating [Big23]. Communication [Dun23]. community [For23, Som23]. Companion [Tal21]. Comparative [Bur22]. Complexity [Hol20]. Compound [Spi23]. compulsions [Ste24]. Computational [Boo23a]. computers [Bel23]. conference [For23, Wid23]. conferences [BRSW23, BS20, Rei23, RBB^{+}20, Som23, Zai23, Kot23]. conferencing [Big23, HLDS20, Wid23]. Congress [Mou23]. consensus [Wes21]. Consequence [Fra20]. conservation [Sch22a, dB20]. Constructing [Ish20]. Contemporary [Dur23]. Contested [Bus22]. context [BD20, HS21, LZ23]. contrasting [Tur23]. conventional [BS20]. Cooley [Tar24]. Corbett [Bro20]. Cornell [Web23]. Correspondence [Bow23b, Bro20]. Corrigendum [Sec21b, Law23a]. cosmology [Bie20]. Costa [Tal21]. Cover [Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21a, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e]. crafting [Kot23]. Craig [Web20]. Creatures [Hem22, Goo20]. credibility [Mat22]. crisis [RBB^{+}20]. critical [SS23, Dee20, Win23a]. Crosland [Smi21a]. crossings [Sch22a]. Csiszar [Pal21]. CT [Amb21b, Kat20, Mid21, Mil21, Mor21a]. cudgelling [Boo23b]. Cueto [Coz20]. Cultivating [LD20]. Cultural [Bee21, Law21a, Law23a, Law23b, Mül24]. Culture [Dol21, Guf21, Jac21, Ros20, Kir20, Hel20]. Cunningham [Jor23]. curated [For23]. Curating [Nic22]. Curies [Med23]. Cyrus [Gro24, Zai23].

D [Bro20, Di 23, Kat20]. Dale [HA20]. Daniel [BR23]. dans [Bow23a]. Darwin [Bow23b, Bow23b, Khr24, Sec21a, Sec21b, Win23a]. Data [Sho22, Bol23, LD20]. David [Ste22, Tal21, Win23a, Sac21]. Death [Web23, Hop22]. debt [Sto22]. decision [Käl22]. decolonization [SS23]. deconstructive [MS21]. Della [Bru22]. demystification [Ame20]. Denis [Sto21]. Descartes [Sch22b]. Design [Des21]. d’Etaples [Ros20]. d’ethnographie [Mau22]. Development [Hol23b, Kan20, Bec23]. dialogues [SS23]. Diarmid [Sec22]. Dick [Guf21]. Dickson [Han20]. digester [Sto21]. digestion [Sch22b]. digital [McG21]. DiMeo [Leh23]. Dinosaur [Bow22]. Dinosaurs [Bow22]. Dipesh [Sim22]. diplomacy [Ada23, Bar23, DUS23, GB23, Med23, TA23, Kri23]. diplomatic [Tur23]. disaster [Hol23c]. Discerning [HA20]. discipline [Joh21]. Disciplines [Jac21]. Discontents [Smi21b]. Discovered [Jor23]. discovery [HS21]. discussion [Bie20]. discussions [Hua21]. disease [Sta22]. dismantling [Bur22]. disorder [Joh21]. Disorders [Hol20]. disputed [Bus22]. Disquiet [Mox21]. distribution [Ish20]. DNA [HS21]. Doctor [Hol22]. Doi [Ran23]. doing [Wil21b]. doppelgängers [Bus22]. Dora [Bru22]. Dorner [Far21]. Dovecote [Ree21]. Down [Kom24]. dread [Sta22]. Dreams [Far20]. Drugs [Sac21]. Dublin [Kir20]. Dubow [GS22]. Duncan [Des20]. Duplicate [Kai22]. Duplicates [Mac22, Bus22, HMB22, Mar22, Nic22]. duplication [Mau22]. during [KK22]. Dutch [tH23]. Dylan [Law21a, Law23a, Law23b]. Dynamics [Boo23a, Dav21].

eagle [SS23]. Earle [LD21]. Early

[Bee21, Nia24, Smi24, Tar22, vD23, Bie20, Bil20, Käl22, Mac22, Mat22, Med23, Sta22]. Earth [Bru22, Mat22, Ros23, Tim20]. Eastern [MB22]. Eat [Smi21b]. Eating [dB20]. Eaton [Zai23]. Eburon [Hol23b]. ecological [Hol23c]. Ecology [Ris20]. Economy [Mid21]. ed [Dee20, Ris20]. Edge [Di 23]. Edhasa [Man24]. Editions [Bow23a]. editors [Bow23b]. Edney [Hay23]. eds [Ant24, Bee21, Boo23a, Bow23b, Bro20, Exl23, Guf21, Jac21, Koh23, Nia24, Sia21, Smi23, Tal21]. Edward [Bie20]. Edwardian [ASW21]. efficiency [Spi20]. Eighteenth [Far21, Jon23, Sto21]. eighteenth-century [Jon23, Sto21]. Einstein [Ste22, Ste22]. Einstein-Marić [Ste22]. Einsteinian [Gor22]. electricity [Wes21]. electronic [McG21]. Elena [Lan22]. elite [Dav21]. Elizabeth [Coz20, Smi21b]. embedding [Led22]. emergency [RBB^{+}20]. emerging [Bek20]. Emma [Mid21]. Empire [Arn23, Kan20, Bil20, KK22, Yal22]. End [Kan20, Lan22, Bow24]. engineer [Dis24]. engineer-like [Dis24]. Engineering [Gro24, Dis24, Web20]. English [Dun22]. Enlightenment [Amb21b, Guf21]. Enough [BR23]. Ensayo [Man24]. enters [Ker20]. entomology [Wor22]. Environment [Bim20, vD23, HA24]. Environmental [Chu23, HA20, HA24]. Environmentalisms [HA21]. Environments [HA21, Ris20]. episode [Dav21]. Epistemology [Dur23]. era [Bow24]. Erika [Koh23, Goo20]. Esguerra [Ant24]. Esterson [Ste22]. estimate [Sto22]. ethnographic [Bus22]. ethological [Nix24]. ethology [Kem23]. Etienne [HA21]. eugenics [Ayl21]. EUR23.00 [Bow23a]. EUR28.00 [Sto23]. EUR36.00 [Hol23b]. EUR85.00 [Nia24]. Eurasian [Ris20]. Europe [Tar22, vD23]. European [Sto21]. Evangelists [LZ24]. Everest [Bim20]. Evolution [Hor23, Tal21, Tur21, Des20, Bow23a]. Evolutionary [Hor23]. Ewing [Kor21]. examining [MB22]. examples [Yal22]. Exhibitions [Win23b]. Expanses [Mey20]. Expedition [Bim20]. Experience [Jon21, HLDS20]. experiences [MB22]. experiment [Bel22, EW23]. Experimental [Dun22]. experimentation [LD20]. Experiments [Hol20, Lan22, Hee23]. expertise [KW21]. Experts [HA20, Kan20]. Exploration [Hol21]. Explorations [McC23]. expressions [Kem23]. Extreme [Bim20, Hol21]. eye [Ber21, Nix24].

facial [Kem23]. failure [Wyn21]. Falling [Tim20]. Falls [Amu23]. family [Ayl21]. famine [LD20]. Fanny [Bru24]. Fantasy [Har21]. Farrah [Koh23]. Fear [Har21]. feasts [Pfe21]. February [Bro20]. Fee [Coz20]. Feeding [LD21]. Felix [Sto23]. fermentation [Sch22b]. fight [Med23]. film [Lon20, tH23]. filogenia [Man24]. Finnegan [Sec22]. Fire [Dun22]. first [Som23, YJ21]. Fisher [Ayl21]. Fiss [Dun23]. Flannery [Tal21]. Fleetwood [Arn23]. Flood [vD23]. Florentino [Man24]. flowers [EW23]. Follow [Jor23]. Fontaine [Di 23]. food [LD20]. forecast [Car21]. forerunner [Khr24]. Forest [Hay21b]. forged [Low21]. Forging [Sle20]. Form [Hor23]. Forms [Des20]. Forum [Ano20g]. Fossil [Bow22, Ree21]. fostering [SS23]. Foundation [Bec23, Amb21a]. Framing [Ano21h, Wil21a]. France [Amb21b]. Frank [Bee21, Boo23b]. Frederick [Bow23b]. Frederico [Amu23]. Freedom [Dua21]. Freitas [Amu23]. French [Amb24, Buc21]. Freud [Hol23a]. Friedrich [HS21]. fringe [Mor21b]. Front [Ano20b, Ano20d, Ano20f, Ano21a, Ano21c, Ano21e, Ano21g, Ano22b, Ano22d, Ano22f, Ano22h, Ano23b, Ano23d, Ano23f]. Frontier [Chu23, Fra21, Ros22]. Frontiers [Boo23a]. fuel [Spi20]. fuel-efficiency [Spi20]. function [Boo23b]. Functional [Kot23]. Funding [Amb21a]. Fungorum [Bol23]. future [Bek20, Big23]. Fyfe [Ran23].

G [Exl23]. Galenic [Spi23]. game [dB20]. Garde [Smi24]. Garden [Hol22]. gastrulation [Hop22]. gathering [BRSW23]. Genealogies [Bru22]. genealogy [McG21]. Genes [McG21]. genetic [Bur22]. Genetics [Hol20, Sha22]. genomics [Low21]. Geographical [Bru22]. Geographien [Sto23]. Geological [Mou23]. Geology [GS21]. Geometry [Wil20, MS21]. Geopolitics [Ant24]. George [Ish20]. German [Bus22]. Germany [Sha22]. Gilded [Sec22]. Gin [Mar20]. Global [Web20, vD23, Mar22]. Globalization [Lor24, MB22, Tek22]. globalizations [Bur22]. Globalizing [Yal22]. go [McG21]. goddess [Som23]. Going [Tim20]. Gonaver [Bla20]. Good [Hes23, BR23]. Gordon [Kot23, Kri23]. Göttingen [Sto23]. Granada [Kem23]. Grant [Boo23a]. gravid [Bel22]. Green [Hol23b, Spi20]. Greenhalgh [Bee21]. Greenwich [Bel23, Ish20, Joh21]. Gresham [Bie20]. Gribenski [Bru24]. Griffin [Mid21]. Gruen [Dee20]. guardians [YJ21]. Guinea [Bus22]. Gunpowder [Wil20].

H [Fra20, Hay23, Win23b]. Haida [Nic22, SS23]. hammer [Mac22]. Handbook [Exl23, Smi23]. Hannah [Koh23]. Hans [Hes23]. hardback [Amb24, Ant24, Arn23, Bee21, BR23, Bow22, Bow23b, Bro20, Bru22, Bud24, Chu23, Di 23, Dol21, Dua21, Dun22, Eic21, Far20, Far21, Far23, Fig21, Fra20, GS21, GS22, Guf21, Har21, Hay21b, Hem22, Hol23a, Hol21, Hol22, HA21, Jac21, Jam21, Jor23, Kan20, Kat20, Koh23, Kor21, Kri23, Lan22, Led21, LD21, Mac20, Mar20, McC23, Mey20, Mid21, Mox21, Nia24, Ols23, Pal21, Ree21, Ros22, Ros20, Sac21, Sec22, Sha22, Sho22, Sia21, Smi20, Smi23, Smi24, Spi23, Ste22, Sto23, Tal21, Tar22, Tur21, Web23, Wil20, vD23]. hardback/paperback [Tar22]. hardcover [Amb21b, Amu23, Boo23a, Des20, Des21, Exl23, Goo20, Gro23, Han20, Hes23, Hor23, Hun23, Lor24, LZ24, Mil21, Mor21a, Ran23, Sle20, Web20]. Harrison [Koh23]. Hartogsohn [Jon21]. Harvard [BR23, Bow22, Gro23, Har21, Pal21]. Harvey [Jor23, McA22]. Haven [Amb21b, Hol22, Jam21, Kat20, Mid21, Mil21, Mor21a]. head [Sec21a, Sec21b]. Health [Coz20, Far21, Dav21]. Heggie [Hol21]. Helen [Mey20]. Henry [EW23, McA22]. hermanos [Man24]. Hertz [Bru24]. Herzberg [Sac21]. Hickman [Hol22]. Hidden [Sac21, Wil21b]. high [Wid23]. high-level [Wid23]. Higher [Hol21]. Hill [Bla20, Chu23]. Himalaya [Arn23]. Hiscock [Mar20]. Historian [Mor21a]. historical [Hee23]. histories [BRSW23]. historiography [MS21]. History [Boo23a, Dun23, Fra20, Gro23, Hem22, Hol21, Hol23b, HA21, Kat20, Koh23, Lan22, Led21, Mey20, Mid21, Mil21, Ran23, Sac21, Sho22, Sia21, Sim22, Smi20, Smi23, Tim20, Tur21, Bel23, Ber21, Bil20, Boo23b, Ker20, Kle20, Kom24, Lon20, Mac22, McA22, Mor21b, MB22, SS23, Bow22, Coz20, Eic21, Hay23, Mül24, Ris20]. Hoffenberg [Win23b]. hole [GB23]. Hong [Tim20, Hua21, Wil21b]. Hopkins [Dua21, Fra20, GS21, Led21, Ols23, Web20, vD23]. hormone [Bec20]. Horticulture [Hol22]. house [Kai22]. household [EW23]. Hudson [Bow22]. humaine [Bow23a]. Human [Eic21, Goo20, Tur21, Bel22, Des20, Tur21]. Humanities [Hab24, Hay21a]. Humanity [Tur21]. Humboldt [Ant24]. Humphry [Des21]. Hunt [Goo20]. Hunter [Bel22]. Hunters [Bow22]. Hutton [Wil20].

IAEA [Ada23]. Ian [Des20].

iatromechanism [Sch22b]. Ibn [Tek22]. Icy [McC23]. idea [HA24]. Ideal [Hay23]. Ido [Jon21]. Iguazu [Amu23]. illusions [Ame20]. im [Sto23]. image [TA23]. images [GB23]. Imagination [Amb24, GS22]. Imagining [vD23]. Impact [Bee21]. Imperial [Ris20, Ano21h, DUS23, Wil21a, Wor22]. impossibility [Car21]. Impossible [Sto22]. Incomparable [Leh23]. India [Ros22, Mil21, YJ21]. Industry [Bud24]. informality [Kot23]. Inheritance [McG21]. Innovation [Guf21, RBB^{+}20]. Innovations [Mey22]. inoculation [Mey22]. Inscriptions [GS21]. insects [Kle20]. Insider [Led21]. Institute [Low21]. Institution [Bla20, Nic22]. institutions [Kai22]. instrument [Ada23]. instruments [Kom24]. Intellectual [Smi23]. interaction [Kot23]. interested [Sto22]. international [Ada23, BRSW23, Low21, Mou23, Som23, dB20, Bec23, Mou23]. internationalism [Rei23]. internationalisms [Zai23]. internationalization [Bol23]. interwar [Ayl21, Lon20]. Intimate [Mid21]. Introduction [Dur23, HMB22, TA23, Hol23c]. Inventing [Dun22]. Invention [Ant24]. investigation [Bel22]. invisible [Mat22]. Ireland [Kir20]. Irina [Man24, Man24]. ISBN [Amb21a, Amb21b, Amb24, Amu23, Ant24, Arn23, Bee21, BR23, Bim20, Bla20, Boo23a, Bow22, Bow23a, Bow23b, Bro20, Bru22, Bru24, Buc21, Bud24, Chu23, Coz20, Dee20, Des20, Des21, Di 23, Dol21, Dua21, Dun22, Dun23, Dur23, Eic21, Exl23, Far20, Far21, Far23, Fig21, Fra20, GS21, GS22, Goo20, Gro23, Gro24, Guf21, Hab24, Han20, Har21, Hay21b, Hay23, Hel20, Hem22, Hes23, Hol23a, Hol20, Hol21, Hol22, Hol23b, HA20, HA21, Hor23, Hun23, Jac21, Jam21, Jon21, Jor23, Kan20, Kat20, Koh23, Kor21, Kri23, Lan22, Law21a, Law23a, Law23b, Led21, Leh23, LD21, Lor24, LZ24, Mac20, Man24, Mar20, McC23, McM20, Mey20, Mid21, Mil21, Mor21a, Mox21, Mül24, Nia24, Ols23, Pal21, Ran23, Ree21, Ris20, Ros22, Ros20]. ISBN [Sac21, Sec22, Sha22, Sho22, Sia21, Sim22, Sle20, Smi20, Smi21b, Smi23, Smi24, Spi23, Ste22, Sto23, Tal21, Tar22, Tar24, Tim20, Tur21, Web20, Web23, Wil20, Win23a, Win23b, vD23]. Islamic [Tek22]. issue [Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21a, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e, HMB22, Mar22]. Italian [Bec20]. Ithaca [Web23]. itself [Low21]. Ivanova [Amb24]. ix [Coz20, HA20].

J [Bow22, Mac20, Ros20, Smi23]. J. [Buc21, Win23a]. Jacobsen [Hol23a]. Jacqueline [Far23, Fra20]. Jahrhundert [Sto23]. Jamaica [Hol23c]. James [Bee21, Bow23b, Tal21, Guf21]. Jamieson [HA20]. Jan [Kle20]. Janet [Bro20]. Jefferson [Di 23]. Jeffrey [LZ24]. Jennifer [Han20, Dun22]. Jersey [Tim20]. Jessica [HA20]. Joan [Kor21]. JoAnne [Web20]. job [Wil21b]. Johan [Hol23b]. John [Bro20, Pie20]. Johns [Dua21, Fra20, GS21, Led21, Ols23, Web20, vD23]. Jones [Exl23, Koh23, Sia21]. Jorge [Ant24]. Joseph [Bar23]. Joshua [Sle20]. Journal [Pal21]. Journals [Ran23]. Julie [Ran23]. Just [Jam21, Wil21b].

Kaalund [McC23]. Kampourakis [Dur23]. Kantorovich [BD20]. Katrine [McC23]. Keith [Kor21]. Keller [Nia24]. Ken [Bro20]. Kerri [Smi20]. Kevin [Dur23]. Kew [Mor21a]. Keynyn [HA20]. Khaldun [Tek22]. Kingston [Hem22]. Knowledge [Dur23, Pal21, Jon23, Led22, Lon20, Mat22, Ant24]. Kong [Tim20, Hua21, Wil21b]. Kostas [Dur23].

Lab [Led21]. laboratories [Bur22].

Laboratory [Dua21, Har21, Wyn21, Led21]. Lachlan [Arn23]. Lady [Leh23]. Landscape [Des21, Mac22, Chu23]. language [Gor22]. late [Bil20, Yal22]. Launius [Kat20]. Law [Kor21]. Lawrence [Koh23, Led21]. Lawrence-Mackey [Koh23]. lay [Bol23]. leadership [YJ21]. leading [Low21]. League [For23]. Lebanon [Bur22]. Lebens [Sto23]. Lecture [Sec22]. Lee [Gor22]. Leeuwenhoek [tH23]. Lefèvre [Ros20]. legacy [Med23]. Leonid [BD20]. Levant [Mey22]. level [Wid23]. Liba [Sia21]. Library [Mar20]. Life [Leh23, Ree21, Wil20, Hop22, Far23]. Light [Hem22]. Lightman [Jac21]. like [Dis24]. linguistic [Gor22]. Linthorst [Hol23b]. Liquidator [McM20]. literary [KK22]. Liu [Mil21]. Livermore [Led21]. Liverpool [Guf21]. Lives [Win23a, Hol23a]. Living [Hem22]. local [Ano21h, Wil21a]. locusts [Wor22]. London [Amb21b, Amb24, Ant24, Bee21, Bow22, Dun22, Dur23, Eic21, Exl23, Fig21, Har21, Hel20, Hol23a, Hol20, Hol21, HA20, Jam21, Jor23, Koh23, Mac20, Mey20, Mil21, Mox21, Ran23, Smi20, Smi24, Tal21, Tim20, Tur21, Wil20, Win23a, Des21, Kem23, Law21a, Law23a, Law23b, Pfe21]. Long [Far21, Gro24, Hay21b]. lore [KW21]. Lorei [Dee20]. Lorraine [Goo20]. Lüders [McC23]. Lukas [Bow22]. Luminous [Hem22]. lunar [Bel23]. Lüttge [Sto23]. lviii [Bro20]. Lydia [vD23].

M [Coz20, Dun22, Gro24, Law21b, Mey20, Mox21, Sec21a, Sec21b]. MA [Amb21a, BR23, Bow22, Far23, Gro23, Gro24, Har21, Jon21, Mül24, Pal21, Ste22]. Mackenzie [Tar24]. Mackey [Koh23]. Macmillan [Exl23]. Magic [Tar22]. Mahoney [Kor21]. maintenance [Kir20]. Making [Bla20, Bow22, KW21, Käl22, Ros20]. malacology [Mar22]. Malcolm [Guf21]. Malinowski [Mar22]. Malthus [Sec21a, Sec21b]. Man [McG21]. management [Dis24]. Managing [Joh21]. Manchester [Kan20]. Manning [Tur21]. Mantle [Bru22]. Mapping [Mat22, Ols23]. maps [Mat22]. Marci [Lor24]. Marco [Hor23]. Marcos [Coz20]. Margarete [Smi24]. Margery [EW23]. Maria [Bud24, Mox21]. Marić [Ste22]. Marie [Nia24]. Mark [Amb21a, Ant24, Tar22]. Market [Sac21]. Marketplace [Bud24]. marriage [Hop22]. Mars [Sle20]. Marsh [Far20]. Martin [Exl23, Koh23, Tim20]. Mary [Ree21]. mask [SS23]. Mass [Sle20]. material [Hee23, Kai22]. Mathematical [Ros20]. Mathematician [Wil20]. Mathematics [Dun23, MS21]. matter [Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21a, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e]. Matthew [HA20, Hay23]. Maurice [Smi21a]. Max [Gor22]. Maynard [Pie20]. McCain [Dur23]. McDougal [Ran23]. McDougal-Waters [Ran23]. McGill [Hem22]. McGill-Queen [Hem22]. McMillan [Bro20]. measurement [Dis24]. Mechanical [Amb21b, Jon23]. Mechanics [Buc21]. mechanisms [Led22]. Media [Sle20, HLDS20]. mediatic [Med23]. mediating [Big23]. Medical [Dav21]. Medicine [Han20, Hol22, KK22, Smi21b, ASW21, Mey22]. Medicines [Far21]. Mediocrity [BR23]. Mediterranean [Spi23]. Melissa [Han20]. memory [tH23]. Mendelian [McG21]. Mendelism [Sha22]. mensuration [Dis24]. Merchants [Far21]. Merpeople [Eic21]. metaphor [Wes21, Bru22]. Metcalf [Ols23]. meteorologists [Wil21b]. method [Bei21]. MI5 [Kor21]. Michael [Bow22, Tal21, Bro20, Hel20, HA20, Jam21]. Michel [Gro23, Hem22]. Michelle [Leh23]. microscope [Bei21]. Microscopical [Bei21]. mid [Spi20]. mid-twentieth [Spi20]. Middle [MB22]. Miescher [HS21]. Migration [Chu23]. Mikkel [Hol23a]. Milam [Goo20]. Milena [HA20, Amb24]. Mileva [Ste22]. Milo [BR23]. Mind [Law21b, Ber21]. Minster [Ree21]. miracles [Ros23]. mirages [Ame20]. Mitton [Far23]. Mobility [Chu23]. Model [Hol20]. Modern [Bee21, Bla20, Hun23, Nia24, Tar22, vD23, Bie20, KW21, Mat22, MS21]. Modernity [Han20]. Mody [Gro24]. Molecular [Gro23]. monsoon [Car21]. Montano [Mox21]. Montreal [Hem22, GB23]. Moon [Kat20]. Morag [Bee21]. Morange [Gro23]. Moretta [Kor21]. Mørk [Ran23]. Morley [Fig21]. Mount [Bim20]. Moxham [Ran23]. Mr [Mor21a]. Multifarious [Mor21a]. Mulvin [Law21a, Law23a, Law23b]. Muqaddimah [Tek22]. Murphy [Web20]. Musée [Mau22]. museumization [SS23]. museums [Boo23b]. Musgrave [Mor21a]. Music [Bru24]. my [Boo23b]. mycology [Bol23]. mysteries [Bol23].

N [Web20]. nacional [Man24]. Nall [Sle20]. Nanna [McC23]. Naomi [HA20]. Narratives [McC23]. NASA [GB23]. Nasim [Mül24]. National [Amb21a, Amu23, Led21]. Nationalizing [Amu23]. Natural [Kle20, Mor21a, Mox21, Tal21, Kom24, Lon20, Mac22, McA22, Ros23]. natural-history [Mac22, McA22]. Naturalization [GS21]. Nature [Amu23, BR23, Fig21, GS21, Goo20, Hel20, Ris20, Tar24, Boo23b, Hol23c, Tal21]. NC [Chu23]. necessary [Led22]. necessity [Car21]. Necropolis [Web23]. Needham [Bar23]. Negotiating [Mou23]. negotiations [Bus22, GB23]. Nelson [Hol20]. Neptune [Har21]. network [Wid23]. networks [KW21, Kai22, Low21]. never [Spi20]. News [Sle20]. Newtonian [Buc21]. Niche [Bec23]. Nicholas [Ris20]. Nicole [Hol20]. nihilists [Khr24]. Nineteenth [Han20, Hay21b, McC23, Pal21, Ros22, Dis24, KK22, Mac22, McA22]. Nineteenth-Century [Han20, Dis24, Mac22, McA22]. NJ [Des20, Dun23, Goo20]. Noah [Ran23]. Noakes [Hun23]. Nobbs [Chu23]. Nobbs-Thiessen [Chu23]. Nobel [Käl22]. noise [Tur23]. normalizing [Bec20]. Norman [Bro20]. norms [Käl22, Mou23]. North [Bla20, Chu23, McC23, Bek20, Hel20]. Northern [Mac20]. Nos [Bow23a]. Novel [Des20]. November [Bro20]. Nuclear [Sho22, YJ21]. nuclein [HS21]. NY [Web23].

object [Bel22]. observation [Fra21, Nix24]. Observatory [Bel23, Joh21, Gil20]. observing [Kle20]. Occult [Hun23]. OCD [Ste24]. Oceans [Mey20]. Old [Amb21b]. Omar [Mül24]. one [Ber21]. online [BS20, HLDS20]. ontology [Bel22]. Oosterhoff [Ros20]. open [Bee21, Koh23, Ran23, KW21]. open-access [Koh23, Ran23]. operationalizing [Nic22]. Oppenheimer [HA20]. Optimizing [Bec20]. order [Joh21]. O’reilly [HA20]. Oreskes [HA20]. Organisms [Hem22]. Organization [Coz20]. Origin [Jac21]. original [HS21]. Origins [Buc21]. Other [Amb21a]. Ottoman [Bil20, KK22, MB22, Sto22, Yal22]. Our [Win23b]. Own [Win23b, Ber21]. Oxford [Kor21, Mar20, Ros20]. ozone [GB23].

P [Wor22]. PA [Bro20, Bud24]. Pacific [Bek20]. Palgrave [Exl23]. pandemic [RBB^{+}20]. Paola [Amb21b]. paperback [Amb21a, Bee21, Bim20, Bla20, Bow22, Bow23a, Buc21, Coz20, Dee20, Di 23, Dua21, Dur23, Fig21, Gro24, Guf21, Hay23, Hel20, Hol20, Hol23b, HA20, HA21, Jon21, Koh23, Law21a, Law23a, Law23b, McM20, Mül24, Pal21, Ris20, Sim22, Smi21b, Ste22, Tar22, Tim20, Tur21, Win23a, vD23]. paperwork [Kle20]. Papin [Sto21]. Paris [Bow23a, Dol21]. Parks [Amu23]. Parliament [Led22]. Parré [Guf21]. Past [Dol21, Big23]. Patients [Hol23a]. Patrick [Tur21]. Paul [Bee21]. Paula [Spi23]. PDF [Koh23, Ran23]. Pearson [Web23]. peces [Man24]. Peculiar [Bla20]. Penser [Bow23a]. People [LD21]. perceptions [Mey22]. Perfection [Tar24]. performance [Ame20]. period [Bie20]. Peter [Mac20, Win23b]. Petty [Kom24]. Pharma [Sac21]. Pharmacy [Spi23]. Philip [Bim20]. Philippe [Di 23]. Philosophy [Boo23a, Mox21, Tek22]. photographs [Bar23, Mau22]. photography [Nix24]. Physical [Gro24, Ame20]. physician [Kle20]. physics [Ber21, Käl22, Hun23]. Physiology [Hol21]. Pickren [Smi23]. Picturing [Bar23]. Pierre [Smi21a]. pig [Low21]. Pipeline [Sch22a]. Pit [Wil20]. Pittsburgh [Bim20, Boo23a, Bro20, Bud24, Han20, Hes23, Hor23, Lor24, LZ24, McC23, Ris20, Sec22, Sle20, Spi23, Win23b]. pivotal [Dav21]. place [Bie20]. plague [Hol23c]. Planetary [Sim22, Bel23]. plant [Kai22]. Plaskett [Mac20]. plots [Tur23]. Podgorny [Man24]. Polar [KW21]. Pole [Hel20]. Policy [Di 23, Wid23, HA20]. political [Wes21]. Politics [Bru24, Hol23b, Lan22, LD21, Pal21, Sha22, Web23, Ste24, dA20]. Pooley [Di 23]. Pop [Hay21b]. population [Bec20]. Portugal [DUS23]. Portuondo [Mox21]. Post [Dol21, Tur23]. post-punk [Tur23]. Post-revolutionary [Dol21]. Postal [Bei21]. Postwar [Di 23]. Potato [LD21]. Power [Ros22, TA23, EW23]. Pp [Amb21a, Amb21b, Amb24, Amu23, Ant24, Arn23, Bee21, BR23, Bim20, Boo23a, Bow22, Bow23a, Bow23b, Bro20, Bru22, Bru24, Buc21, Bud24, Chu23, Coz20, Dee20, Des20, Des21, Di 23, Dol21, Dua21, Dun22, Dun23, Dur23, Eic21, Exl23, Far20, Far21, Far23, Fig21, Fra20, GS21, Goo20, Gro23, Gro24, Guf21, Hab24, Han20, Har21, Hay21b, Hay23, Hel20, Hem22, Hes23, Hol23a, Hol20, Hol21, Hol22, Hol23b, HA20, HA21, Hor23, Hun23, Jac21, Jam21, Jon21, Jor23, Kan20, Kat20, Koh23, Kor21, Kri23, Lan22, Law21a, Law23a, Law23b, Led21, Leh23, LD21, Lor24, LZ24, Mac20, Man24, Mar20, McC23, McM20, Mey20, Mid21, Mil21, Mor21a, Mox21, Mül24, Nia24, Ols23, Pal21, Ran23, Ree21, Ris20, Ros22, Ros20, Sec22, Sha22]. Pp [Sho22, Sia21, Sim22, Sle20, Smi20, Smi21b, Smi23, Smi24, Spi23, Ste22, Sto23, Tal21, Tar22, Tar24, Tim20, Tur21, Web20, Web23, Win23a, Win23b, vD23, GS22, Wil20]. practicalities [McA22]. practice [Käl22]. Practices [HA20]. Pratik [GS21]. precarity [Low21]. presence [Big23]. present [Big23, Chu23, GS22]. Presidential [Boo23b]. Press [Amb21a, Amb21b, Amu23, Arn23, Bee21, BR23, Bim20, Bla20, Boo23a, Bow22, Bow23b, Bro20, Bru22, Bru24, Buc21, Bud24, Chu23, Coz20, Dee20, Des20, Di 23, Dol21, Dua21, Dun22, Dun23, Far20, Far21, Far23, Fig21, Fra20, GS21, GS22, Goo20, Gro23, Gro24, Guf21, Hab24, Han20, Har21, Hay23, Hem22, Hes23, Hol20, Hol21, Hol22, HA20, HA21, Hor23, Hun23, Jam21, Jon21, Kan20, Kat20, Koh23, Kor21, Kri23, Lan22, Law21a, Law23a, Law23b, Led21, Leh23, LD21, Lor24, LZ24, Mac20, McC23, Mid21, Mil21, Mor21a, Mox21, Mül24, Ols23, Pal21, Ran23, Ree21, Ris20, Ros22, Ros20, Sac21, Sec22, Sha22, Sho22, Sia21, Sim22, Sle20, Smi21b, Smi23, Spi23, Ste22, Tal21, Tar22, Tar24, Tur21, Web20, Web23, Win23b]. Press [vD23]. primate [Kem23]. Princeton [Des20, Goo20]. Print [Ros20]. Priorities [Dav21]. prizes [Käl22]. problem [Ano21h, Wil21a]. Production [Hem22, KK22, Lon20]. productivity [Kot23]. Program [Hab24]. Programming [BD20]. Project [Bow23b]. projecting [Som23]. Prophet [Tal21]. proteins [dB20]. Protocol [GB23]. provide [Sto22]. providence [Ros23]. Proxies [Law21a, Law23a, Law23b]. Psychedelic [Jon21]. Psychiatric [Hol20]. Psychiatry [Bla20, Wyn21]. Psychics [Hun23]. psychology [Ste24, Smi23]. Psychotechnics [Smi24]. Public [Amb21a, Di 23, Win23b, Boo23b, Sto22]. Publishing [McM20, Ran23, McG21]. Pugwash [Zai23]. punk [Tur23]. Purchase [Dol21].

Queen [Hem22]. quotation [Hop22].

R [Lor24, Mac20]. R. [Ayl21]. Race [Sha22, Tar24, Kat20]. Radder [Hes23]. Radiation [LZ24]. radio [Pie20, Tur23]. radiocarbon [Ker20]. Radium [Bud24, Kir20]. railways [Gil20]. Rampling [Dun22]. Ramsey [Boo23a]. Ranelagh [Leh23]. rattles [Nic22]. raw [Bel23]. Re [MB22]. Re-examining [MB22]. Reaching [Kat20]. Reaktion [Des21, Eic21, Hel20, Hol23a, Mey20, Smi20, Tur21, Win23a]. Real [Ste22]. Rebecca [Koh23, LD21]. Rebel [Wil20]. rebranding [dB20]. reception [Khr24]. Reconstructing [Hes23]. reductions [Bel23]. Rees [Tur21]. reflections [Big23]. Reform [For23]. Regime [Amb21b]. regret [Fra21]. relativity [Gor22]. Religion [Tar22, Yal22]. Remaking [Arn23, tH23]. Remedies [Spi23]. remedy [Sta22]. Renaissance [Tar24]. Rentetzi [Bud24]. representation [Hee23]. Representing [Fig21, Tur23]. Repton [Des21]. republican [Bil20]. Research [Dav21, Kot23, Low21, Hol23b, HLDS20]. responsibility [Pie20]. Restricted [Sho22]. resuscitation [Pfe21]. rethinking [RBB^{+}20]. Retrospectives [Bil20]. Review [Amb21a, Amb21b, Amb24, Amu23, Ant24, Arn23, BR23, Bim20, Bla20, Boo23a, Bow23a, Bow23b, Bro20, Bru22, Bru24, Buc21, Bud24, Chu23, Coz20, Dee20, Des20, Des21, Dol21, Dun22, Dun23, Dur23, Eic21, Exl23, Far20, Far21, Far23, Fra20, GS21, GS22, Goo20, Gro23, Gro24, Hab24, Han20, Har21, Hay21b, Hay23, Hel20, Hem22, Hes23, Hol23a, Hol20, Hol21, Hol22, Hol23b, HA20, Hor23, Hun23, Jac21, Jam21, Jon21, Jor23, Kan20, Kat20, Koh23, Kor21, Kri23, Lan22, Law21a, Law23a, Law23b, Led21, Leh23, LD21, Lor24, LZ24, Mac20, Mar20, McC23, McM20, Mey20, Mid21, Mil21, Mor21a, Mox21, Mül24, Nia24, Ols23, Ran23, Ree21, Ris20, Ros22, Ros20, Sac21, Sec22, Sha22, Sho22, Sia21, Sim22, Sle20, Smi20, Smi21b, Smi23]. Review [Smi24, Spi23, Sto23, Tar22, Tar24, Tim20, Web20, Web23, Wil20, Win23a, Win23b]. Reviews [Bee21, Bow22, Di 23, Dua21, Fig21, Guf21, HA21, Man24, Pal21, Ste22, Tal21, Tur21, vD23]. revisited [Law21b]. Revolution [Bow22, Des21, Lan22, Gro23]. Revolutionary [Wes21, Dol21]. Revolutions [Sec21a, Sec21b]. revolving [Hua21]. Rewriting [Bow22]. Richard [Hun23, Ros20]. Rieppel [Bow22]. Rise [Bru24, Win23b]. Risk [Fra20]. rivalries [DUS23]. river [Dis24]. road [Bec23]. Robert [Leh23, Bee21, Law21b, Sec21a, Sec21b, Smi23]. Roger [Kat20]. role [Boo23b, Low21, Spi20]. Roof [Arn23]. Roos [Nia24]. Rosario [Man24]. Roslin [Low21]. Røstvik [Ran23]. Routledge [Amb24, Ant24, Dur23, Jac21, Jor23, Smi24]. Royal [Fig21, Ran23, Bel23, Gil20, Kir20]. Rozwadowski [Mey20]. Rubin [Far23]. Rule [Kor21]. Rules [Web20]. Rurik [Bek20]. Russel [Tal21]. Russia [Khr24]. Russian [Ris20]. Rutgers [Dun23].

S [BR23, HA21, Mac20]. Sabine [Kan20]. Saccardo [Bol23]. Sadie [Koh23]. safe [Bei21]. sailor [Jon23]. Sally [Han20]. Satellite [GB23]. Saul [GS22]. scholarship [RBB^{+}20]. School [HA24]. Science [Amb21a, Amb21b, Amb24, Bee21, Boo23a, Bru24, Bud24, Di 23, Exl23, Fig21, Har21, Hay21b, Hem22, Hes23, Hol22, Hol23b, Hor23, Käl22, Koh23, Kri23, McC23, Ros22, Sec22, Sia21, Smi21b, Smi24, Tar22, Tim20, TA23, Win23b, dA20, Ada23, Ano21h, Bar23, Bec20, Bek20, Big23, Bil20, Boo23b, For23, Gil20, Hee23, Led22, Mor21b, Mou23, MB22, SS23, Ste24, Tek22, Wid23, Wil21a, Yal22, tH23, Arn23, Bec23, Bim20, Boo23a, Dua21, Dur23, Kan20, Sia21, Win23b]. Sciences [Hab24, Hun23, HA24]. Scientific [Bow22, Dur23, GS22, Hol23b, HA20, McM20, Pal21, Pie20, Ran23, Tim20, Wil20, BRSW23, Käl22, Kir20, Kot23, Mey22, Wyn21, ASW21, Lan22]. scientist [Zai23]. Scientists [Gro24, Sec22]. Scientizing [HA24]. Sclavo [Sta22]. Scribner [Eic21]. scrutiny [Led22]. Sea [Har21, Bek20, KW21]. seamanship [Jon23]. seasonal [Car21]. Secord [Bow23b]. Secrecy [Sho22]. Section [Fra20]. security [LD20]. Seduced [Bud24]. Seeing [Ber21]. Selection [Tal21]. Semi [Web23]. Semi-colonial [Web23]. sensation [Bel22]. sense [Bel22]. Sergei [McM20]. serial [Som23]. sermons [Pfe21]. serum [Sta22]. serve [Som23]. service [Khr24]. Set [Jon21]. Setting [Jon21, Web20]. Seuil [Bow23a]. seventeenth [Pfe21]. seventeenth-century [Pfe21]. Sex [For23]. Sexual [For23]. Shanghai [Tim20]. Shank [Buc21]. Shaped [McC23, Mor21a]. shaping [MS21]. sharing [HLDS20]. Sharpe [Ree21]. Sherwood [Boo23b]. Shiach [Bee21]. Shindell [HA20]. shock [Wes21]. Short [Kat20, Ker20]. Showcasing [Ada23]. Shuttleworth [Han20]. Siam [Web23]. Siam/ [Web23]. Silk [Far20]. similarity [Nic22]. Simon [Far23, Mar20]. Simpson [Ros22]. since [Exl23, Web20]. Singapore [McM20, Tim20]. Sister [Leh23]. sixth [Tek22]. Sleigh [Tur21]. slides [Bei21]. smallpox [Mey22]. Smith [Tal21, Pie20]. Smithsonian [Nic22]. Sociability [Kir20]. Social [Di 23, HLDS20, Kot23, Pie20, Amb21a, Sha22]. Society [Fig21, Hol23b, Ran23, BR23, Di 23, Hes23, Kir20, Pfe21]. softcover [Man24]. Solly [HA24]. Solovey [Amb21a]. Some [Boo23b]. Soul [Dua21]. Sound [Hab24]. sounds [Dis24]. Sourcebook [Koh23]. South [GS22, Car21]. Sovereign [Web23]. Soviet [Ris20, Smi24]. Space [Kat20, Ros22]. Spain [Spi23]. Spanish [Mox21]. speaking [Ber21]. specimens [Mau22]. Spectacle [Bow22]. spore [Bol23]. Spuren [Sto23]. Squares [Gro24]. stacked [Tur23]. stakeholdership [Sch22a]. Stammers [Dol21]. Standard [Web20]. standardization [Mou23]. Standing [Law21a, Law23a, Law23b]. Star [Mac20]. state [Spi20]. States [Di 23, Sho22]. stations [Fra21]. statistics [Kom24]. Sternberg [Smi23]. Steven [Amb24, Kle20]. stimulants [ASW21]. Stoppani [LZ23]. storms [Hua21]. Story [Ste22, McM20]. strategies [DUS23]. Stratifying [Jon23]. Struggle [Dua21]. Studies [Dee20]. study [Kir20, Kle20]. Super [Wid23]. surrounding [Bus22]. Surroundings [HA21]. Survey [Kom24]. Sustainability [Hay21a]. Swammerdam [Kle20]. Sweden [Bec23]. Sylloge [Bol23]. Symbols [Hay21b]. system [Ish20, Sch22a, Tur21]. systems [Mar22].

T [Tal21]. Taipei [Tim20]. Talmey [Gor22]. Tamborini [Hor23]. Tarter [Led21]. Taub [Sia21]. taxonomy [Bol23]. Taylor [Boo23b]. Tea [Mil21]. teaching [Ber21, Hee23]. Technical [Rei23]. technique [Rei23]. Technology [Fra20, Hes23, LZ24, Smi24, Spi20, TA23]. Teicher [Sha22]. telegraphic [Ish20]. teleology [Sch22b]. television [Kem23, Pie20]. term [Dis24]. Terms [Dee20]. territorial [Kom24]. Thames [Bow22]. their [Ame20, Wil21b]. Theodore [Coz20]. Theology [Tal21]. Theory [Ros23, Spi20]. therapies [Bec20]. Therapy [LZ24, Sta22]. Thiessen [Chu23]. Thinking [Hab24]. Thomas [Ros22, Ros23].

Thorogood [Mar20]. Thurner [Ant24]. Tibet [Fra21]. ticks [Hol23c]. time [Ish20]. Times [Han20]. Tkaczyk [Hab24]. Toby [Mor21a]. Tokyo [Tim20]. Tolerance [BR23]. Tom [Des21, Dol21, Ree21]. tools [GB23]. Toronto [Mac20]. trace [Ame20]. Trais [Web23]. Trans [Sch22a]. Trans-Alaska [Sch22a]. Transformations [Hee23]. Transformed [Bud24]. transits [Gil20]. translated [HS21]. translation [Khr24, Tek22]. transmission [Bei21]. transnational [Med23]. Travel [McC23, ASW21]. treatments [Med23]. Tremoring [Gil20]. Trip [Jon21]. Trocadéro [Mau22]. Tropical [Chu23, ASW21]. troubles [McA22]. Truth [Hay21b]. Tuning [Bru24]. Turkey [Bil20]. Turn [LZ24]. Turnhout [Nia24]. Tuscaloosa [Tal21]. Twentieth [Bee21, Hor23, Jon21, Kir20, Spi20, Sta22, Wyn21]. Twentieth-Century [Hor23, Kir20, Sta22, Wyn21]. Two [BS20, Hua21]. Tycoons [Bow22]. Tyndall [Bro20]. typhoon [Hua21].

UCL [Bee21, Koh23, Ran23]. UK [Led22]. Ultimatum [DUS23]. Uncertainty [LZ24]. Underground [Tim20]. Understanding [Amb24]. understandings [ASW21]. Unfriendly [YJ21]. Union [Smi24]. United [Di 23, Sho22]. University [Amb21b, Amu23, Arn23, BR23, Bim20, Bla20, Boo23a, Bow22, Bow23b, Bro20, Bru22, Bru24, Buc21, Bud24, Chu23, Coz20, Dee20, Des20, Di 23, Dol21, Dua21, Dun22, Dun23, Far20, Far21, Fig21, Fra20, GS21, GS22, Goo20, Gro23, Guf21, Hab24, Han20, Har21, Hay23, Hem22, Hes23, Hol20, Hol21, Hol22, HA20, HA21, Hor23, Hun23, Jam21, Kan20, Kat20, Kor21, Kri23, Lan22, Led21, Leh23, LD21, Lor24, LZ24, Mac20, Mar20, McC23, Mid21, Mil21, Mor21a, Mox21, Ols23, Pal21, Ris20, Ros22, Ros20, Sac21, Sec22, Sha22, Sho22, Sia21, Sim22, Sle20, Smi21b, Smi23, Spi23, Tal21, Tar22, Tar24, Tur21, Web20, Web23, Win23b, vD23]. Unravelled [Far20]. Urge [Smi21b]. Uses [Bil20]. using [HLDS20]. USSR [BD20]. uterus [Bel22]. Utrecht [Hol23b]. Uvarov [Wor22].

V [BD20]. value [Mar22]. Vanessa [Hol21]. Vast [Mey20]. Vaughn [Eic21]. Vera [Nia24, Far23]. Verlag [Sto23]. Veronica [Bru22]. very [Dis24]. Vestiges [Khr24]. Victoria [Hab24]. Victorian [Mid21, ASW21, Gil20, Jac21]. vienen [Man24]. Visible’ [Ste24]. visions [Mat22]. Visual [Mau22, Mül24, Ada23, DUS23, GB23, Med23, TA23]. visualizations [GB23]. Viterbo [Man24]. Vöhringer [Smi24]. Voice [Sec22]. vol [Bro20, Sia21]. Voltaire [Buc21]. volume [Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21a, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e, Bow23b]. voluntarism [Ros23]. Vos [Spi23]. Voyaging [Bek20].

W [Bim20, Mül24]. Waddell [Tar22]. Wade [Smi23]. Wales [Mor21b]. Walking [Smi20]. Wallace [Tal21]. Wallis [Han20]. Wallstein [Sto23]. Wals [Sto23]. Wanderers [Smi20]. War [Dua21, Goo20, Kor21, Kri23, Mil21, Lan22]. Wardhaugh [Wil20]. wartime [Bar23]. was [Spi20]. Watching [Nix24]. Waters [Ran23]. Watt [Guf21]. Wazeck [HA20]. weasels [Hol23c]. weather [Ano21h, Car21, Fra21, Wil21a]. Wellerstein [Sho22]. Wendy [Bla20]. Wheat [Lor24]. Wheeler [Jam21]. White [Sac21]. Who [Mor21a]. Whorf [Gor22]. Wife [Ste22]. wildlife [dB20]. William [Jor23, Wil20, Bel22, GS22, McA22]. Williams [Smi21b]. Williamson [Des21].

Wills [Koh23]. Wilson [Ber21]. Wimborne [Ree21]. Winner [Mid21]. winter [Bel23]. Wisdom [Nia24]. Wolf [Exl23, Fra20]. Wolfe [Dua21]. Womack [LZ24]. Woman [Ree21]. Women [Exl23, Smi20, Koh23]. Work [Law21a, Law23a, Law23b, Mou23]. works [Mau22]. World [Arn23, Bru24, Coz20, Far20, McM20, Ols23, Tim20, McA22, Wid23, For23, Mor21a]. Wragge [Fig21]. Wragge-Morley [Fig21].

x [Dua21, Far23]. xi [Bud24, Tim20]. xii [Jor23]. xiii [Sec22, Tur21]. xiv [Far20, Hol22, Sha22]. xix [Sia21]. xvi [Mox21, Ros22]. xvii [Bim20, Hem22]. xx [Mac20]. xxviii [Koh23].

Yale [Amb21b, Hol22, Jam21, Kat20, Mid21, Mil21, Mor21a]. Yates [Web20]. years [Boo23b]. York [Hay21b, Jac21]. Young [Law21b, Sec21a, Sec21b].

Zachary [Far21]. Zon [Jac21]. Zone [Hay21b]. Zoo [Kem23]. Zuckerman [HA24].

References

Adamson:2023:SIA

[Ada23] Matthew Adamson. Showcasing the international atom: the IAEA Bulletin as a visual science diplomacy instrument, 1958– 1962. British Journal for the History of Science, 56:205–223, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/showcasing-the-
international-atom-the-iaea-

bulletin-as-a-visual-
science-diplomacy-
instrument-19581962/
74ADFFE477CDA3169B3141A97A1A290E.

Ambler:2021:BRM

[Amb21a] Katherine Ambler. Book review: Mark Solovey, Social Science for What? Battles over Public Funding for the ‘Other Sciences’ at the National Science Foundation. Cambridge, MA: MIT Press, 2020. Pp. 398. ISBN: 978-0-262-53905-

0. $50.00 (paperback). British Journal for the History of Science, 54(1):113–114, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
mark-solovey-social-science-
for-what-battles-over-
public-funding-for-the-
other-sciences-at-the-
national-science-foundation-
cambridge-ma-mit-press-2020-
pp-398-isbn-9780262539050-
5000-paperback/151265806FABF2BEA3F69B0DA4EA3EA3.

Amborg:2021:BRP

[Amb21b] Jens Amborg. Book review: Paola Bertucci, Artisanal Enlightenment: Science and the Mechanical Arts in Old Regime France. New Haven, CT and London: Yale University Press, 2017. Pp. 312. ISBN 978-0-300-22741-

3. $40.00 (hardcover). British Journal for the History of Science, 54(1):130–131, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
paola-bertucci-artisanal-
enlightenment-science-and-
the-mechanical-arts-in-old-
regime-france-new-haven-ct-
and-london-yale-university-
press-2017-pp-312-isbn-
9780300227413-4000-hardcover/
0A5EBBD2A86ACD736BCBB8DFC9F37698.

Ambrosio:2024:BRM

[Amb24] Chiara Ambrosio. Book review:

Milena Ivanova and Steven French, The Aesthetics of Science: Beauty, Imagination and Understanding, London: Routledge, 2022. Pp. 224. ISBN 978-1-032-33718-0. £110.00 (hardback). British Journal for the History of Science, 57(1): 125–127, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
milena-ivanova-and-steven-
french-the-aesthetics-of-
science-beauty-imagination-
and-understanding-london-
routledge-2022-pp-224-isbn-
9781032337180-11000-
hardback/47B7F4A2AAC9EB4674BF4D6AA0EE928E.

Amery:2020:ATI

[Ame20] Fiona Amery. ‘An attempt to trace illusions to their physical causes’: atmospheric mirages and the performance of their demystification in the 1820s and 1830s. British Journal for the History of Science, 53(4):443–467, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/an-attempt-to-trace-
                                        illusions-to-their-physical-
                                        causes-atmospheric-mirages-
                                        and-the-performance-of-
                                        their-demystification-in-
                                        the-1820s-and-1830s/
                                        B6987E17A024924D6939138D6AA0A01A.

Amuchastegui:2023:BRF

[Amu23] Maria Amuchastegui. Book review: Frederico Freitas, Nationalizing Nature: Iguazu Falls and National Parks at the Brazil–Argentina Border Cambridge: Cambridge University Press, 2021. Pp. 333. ISBN 978-1-108-84483-3. $103.00 (hardcover). British Journal for the History of Science, 56(4):579–582, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/frederico-freitas-
                                        nationalizing-nature-iguazu-
                                        falls-and-national-parks-at-
                                        the-brazilargentina-border-
                                        cambridge-cambridge-
                                        university-press-2021-pp-
                                        333-isbn-9781108844833-10300-
                                        hardcover/6FD4245159AA4B561FE5645F62920881.

Anonymous:2020:BVIb

[Ano20a] Anonymous. BJH volume 53 issue 1 cover and back matter. British Journal for the History of Science, 53(1):b1–b4, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-issue-
1-cover-and-back-matter/
0058932D8AAB0E578F856ED5A199303F.

Anonymous:2020:BVIa

[Ano20b] Anonymous. BJH volume 53 issue 1 cover and front matter. British Journal for the History of Science, 53(1):f1–f2, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-issue-
1-cover-and-front-matter/
F40087CC5173FE67ACD0BD426C9E69F2.

Anonymous:2020:BVId

[Ano20c] Anonymous. BJH volume 53 issue 2 cover and back matter. British Journal for the History of Science, 53(2):b1–b4, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-issue-
2-cover-and-back-matter/
6C9C4EFEF420657B595064B14477FD8C.

Anonymous:2020:BVIc

[Ano20d] Anonymous. BJH volume 53 issue 2 cover and front matter. British Journal for the History of Science, 53(2):f1–f2, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-issue-
2-cover-and-front-matter/
34AE32A46E4D651DCBE984A5EC825532.

Anonymous:2020:BVIf

[Ano20e] Anonymous. BJH volume 53 issue 4 cover and back matter. British Journal for the History of Science, 53(4):b1–b4, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-issue-
4-cover-and-back-matter/
407CA57FCF97156253F3606256A9CAB4.

Anonymous:2020:BVIe

[Ano20f] Anonymous. BJH volume 53 issue 4 cover and front matter. British Journal for the History of Science, 53(4):f1–f2, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-53-issue-
4-cover-and-front-matter/
6923463CF99F5B2E58D220DEFFF52B44.

Anonymous:2020:F

[Ano20g] Anonymous. Forum. British Journal for the History of Science, 53(1):??, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
scientific-broadcasting-as-
a-social-responsibility-
john-maynard-smith-on-radio-
and-television-in-the-1960s-
and-1970s/
6F06499A7EDA8C0282E56DB79BA95F9A.

Anonymous:2021:BVIa

[Ano21a] Anonymous. BJH volume 54 issue 1 cover and front matter. British Journal for the History of Science, 54(1):f1–f2, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-54-issue-
1-cover-and-front-matter/
455321967C798D30144F23675BD3875F.

Anonymous:2021:BVIc

[Ano21b] Anonymous. BJH volume 54 issue 2 cover and back matter. British Journal for the History of Science, 54(2):b1–b4, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/

                                        article/bjh-volume-54-issue-
                                        2-cover-and-back-matter/
                                        910735495A9A03ECA1D195554274BD34.

Anonymous:2021:BVIb

[Ano21c] Anonymous. BJH volume 54 issue 2 cover and front matter. British Journal for the History of Science, 54(2):f1–f2, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-54-issue-
                                        2-cover-and-front-matter/
                                        4B348B7B1B26AA55D60263B31D4ECC71.

Anonymous:2021:BVIe

[Ano21d] Anonymous. BJH volume 54 issue 3 cover and back matter. British Journal for the History of Science, 54(3):b1–b4, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-54-issue-
                                        3-cover-and-back-matter/
                                        5EF77664B0EE32FC1A801AFA6748EA7D.

Anonymous:2021:BVId

[Ano21e] Anonymous. BJH volume 54 issue 3 cover and front matter. British Journal for the History of Science, 54(3):f1–f2, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-

for-the-history-of-science/
article/bjh-volume-54-issue-
3-cover-and-front-matter/
B46886B9FCB9DE5B2A4254019517AE56.

Anonymous:2021:BVIg

[Ano21f] Anonymous. BJH volume 54 issue 4 cover and back matter. British Journal for the History of Science, 54(4):b1–b5, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-54-issue-
4-cover-and-back-matter/
E7FE61616D71B77A245A4AB3E55F52DC.

Anonymous:2021:BVIf

[Ano21g] Anonymous. BJH volume 54 issue 4 cover and front matter. British Journal for the History of Science, 54(4):f1–f2, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-54-issue-
4-cover-and-front-matter/
F1A37074075A0BDCC17B8A3449185A9E.

Anonymous:2021:FAA

[Ano21h] Anonymous. Framing Asian atmospheres: imperial weather science and the problem of the local c. 1880–1950. British Journal for the History of Science, 54(3), September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/ancient-lore-with-
                                        modern-appliances-networks-
                                        expertise-and-the-making-of-
                                        the-open-polar-sea-18511853/
                                        9FB26207EDAE5DBD31EE96D5EC829B0B.

Anonymous:2022:BVIb

[Ano22a] Anonymous. BJH volume 55 issue 1 cover and back matter. British Journal for the History of Science, 55(1):b1–b3, March 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-55-issue-
                                        1-cover-and-back-matter/
                                        57A59E54EC0BF2B8AC08947ACC86C761.

Anonymous:2022:BVIa

[Ano22b] Anonymous. BJH volume 55 issue 1 cover and front matter. British Journal for the History of Science, 55(1):f1–f2, March 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-55-issue-
                                        1-cover-and-front-matter/
                                        E5D8287DB2F729BB84764062C2A7602C.

Anonymous:2022:BVId

[Ano22c] Anonymous. BJH volume 55 issue 2 cover and back matter. British Journal for the History of Science, 55(2):b1–b4, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-55-issue-
2-cover-and-back-matter/
7C6A3E61393F8710651FBF50B01D318F.

Anonymous:2022:BVIc

[Ano22d] Anonymous. BJH volume 55 issue 2 cover and front matter. British Journal for the History of Science, 55(2):f1–f2, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-55-issue-
2-cover-and-front-matter/
39C5A56421E11409DB89DA16C25C041A.

Anonymous:2022:BVIf

[Ano22e] Anonymous. BJH volume 55 issue 3 cover and back matter. British Journal for the History of Science, 55(3):b1–b4, September 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-55-issue-
3-cover-and-back-matter/
86A1733562C5BAE49FA2AF8D42F5A621.

Anonymous:2022:BVIe

[Ano22f] Anonymous. BJH volume 55 issue 3 cover and front matter. British Journal for the History of Science, 55(3):f1–f2, September 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-55-issue-
                                        3-cover-and-front-matter/
                                        6AF6A22729353181A0AD498CA9E0FF66.

Anonymous:2022:BVIh

[Ano22g] Anonymous. BJH volume 55 issue 4 cover and back matter. British Journal for the History of Science, 55(4):b1–b4, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-55-issue-
                                        4-cover-and-back-matter/
                                        0E65CD2BA9FE6AB768DEBD818C0B46F3.

Anonymous:2022:BVIg

[Ano22h] Anonymous. BJH volume 55 issue 4 cover and front matter. British Journal for the History of Science, 55(4):f1–f2, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/bjh-volume-55-issue-
                                        4-cover-and-front-matter/
                                        01FB4E52063F5DDF376949B3FCB3F4D7.

Anonymous:2023:BVIb

[Ano23a] Anonymous. BJH volume 56 issue 1 cover and back matter. British Journal for the History of Science, 56(1):b1–b2, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-56-issue-
1-cover-and-back-matter/
95DDB64D05811F6FF802F15B9074E695.

Anonymous:2023:BVIa

[Ano23b] Anonymous. BJH volume 56 issue 1 cover and front matter. British Journal for the History of Science, 56(1):f1–f2, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-56-issue-
1-cover-and-front-matter/
A9A421918B98C972F43EF9ECD087AD89.

Anonymous:2023:BVId

[Ano23c] Anonymous. BJH volume 56 issue 2 cover and back matter. British Journal for the History of Science, 56:b1–b4, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
bjh-volume-56-issue-2-cover-
and-back-matter/
1BC8E28A19A51FDDCCF91FC78AF9D0B8.

Anonymous:2023:BVIc

[Ano23d] Anonymous. BJH volume 56 issue 2 cover and front matter. British Journal for the History of Science, 56:f1–f2, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
bjh-volume-56-issue-2-cover-
and-front-matter/
2582A3954C0F09A1B82A2A3255621B35.

Anonymous:2023:BVIf

[Ano23e] Anonymous. BJH volume 56 issue 3 cover and back matter. British Journal for the History of Science, 56(3):b1–b4, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-56-issue-
3-cover-and-back-matter/
491FE04838AED2E13469EE5413B95E93.

Anonymous:2023:BVIe

[Ano23f] Anonymous. BJH volume 56 issue 3 cover and front matter. British Journal for the History of Science, 56(3):f1–f2, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/bjh-volume-56-issue-
3-cover-and-front-matter/
C012C36C3F76BF0FC4E42C25EB119AFF.

Anthony:2024:BRM

[Ant24] Patrick Anthony. Book review: Mark Thurner and Jorge Cañizares–Esguerra (eds), The Invention of Humboldt: On the Geopolitics of Knowledge, London: Routledge, 2023. Pp. 342. ISBN 978-1-032-13916-6. £96.00 (hardback). British Journal for the History of Science, 57(1):134–136, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/mark-thurner-and-
jorge-canizaresesguerra-eds-
the-invention-of-humboldt-
on-the-geopolitics-of-
knowledge-london-routledge-
2023-pp-342-isbn-
9781032139166-9600-hardback/
5EC7C5A9A8907461F7A682844980BEF7.

Arnold:2023:BRL

[Arn23] Katherine Arnold. Book review:

Lachlan Fleetwood, Science on the Roof of the World: Empire and the Remaking of the Himalaya. Cambridge: Cambridge University Press, 2022. Pp. 294. ISBN 978-1-009-12311-2. £75.00 (hardback). British Journal for the History of Science, 56(4):582–583, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/lachlan-fleetwood-
science-on-the-roof-of-the-

                                        world-empire-and-the-
                                        remaking-of-the-himalaya-
                                        cambridge-cambridge-
                                        university-press-2022-pp-
                                        294-isbn-9781009123112-7500-
                                        hardback/EF437D1372EB122CC05BFD9CBE705A76.

Armston-Sheret:2021:ATM

[ASW21] Edward Armston-Sheret and Kim Walker. Is alcohol a tropical medicine? Scientific understandings of climate, stimulants and bodies in Victorian and Edwardian tropical travel. British Journal for the History of Science, 54 (4):465–484, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/is-alcohol-
                                        a-tropical-medicine-
                                        scientific-understandings-
                                        of-climate-stimulants-and-
                                        bodies-in-victorian-and-
                                        edwardian-tropical-travel/
                                        0B3E72F2CE3EC4E56D609F62C3546CC4.

Aylward:2021:RFE

[Ayl21] Alex Aylward. R. A. Fisher, eugenics, and the campaign for family allowances in interwar Britain. British Journal for the History of Science, 54(4):485–505, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/ra-fisher-eugenics-
                                        and-the-campaign-for-family-
                                        allowances-in-interwar-

britain/
BD28704590AC49EAD4B6FC8F4471C455.

Barrett:2023:PCS

[Bar23] Gordon Barrett. Picturing Chinese science: wartime photographs in Joseph Needham’s science diplomacy. British Journal for the History of Science, 56:185–203, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/picturing-
chinese-science-wartime-
photographs-in-joseph-
needhams-science-diplomacy/
27697B43FABBBBFCC089FBB4F76F427E.

Boldyrev:2020:PUL

[BD20] Ivan Boldyrev and Till Düppel.

Programming the USSR: Leonid V. Kantorovich in context. British Journal for the History of Science, 53(2):255–278, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
programming-the-ussr-leonid-
v-kantorovich-in-context/
4BF0F0D89079DD94AF595EA25A991299.

Beccalossi:2020:ONP

[Bec20] Chiara Beccalossi. Optimizing and normalizing the population through hormone therapies in Italian science, c. 1926–1950. British Journal for the History of Science, 53(1):67–88, March 2020.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
optimizing-and-normalizing-
the-population-through-
hormone-therapies-in-
italian-science-c19261950/
766EBC9D01796500ADA4EA4CE1AFFEA7.

Beckman:2023:NDI

[Bec23] Jenny Beckman. Niche development: the International Foundation for Science and the road to Sweden. British Journal for the History of Science, 56(4): 553–566, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
niche-development-the-
international-foundation-
for-science-and-the-road-to-
sweden/
F6FA6A9AE66F24725778556423EE284A.

Beeston:2021:BRR

[Bee21] Erin Beeston. Book reviews: Robert Bud, Paul Greenhalgh, Frank James and Morag Shiach (eds.), Being Modern: The Cultural Impact of Science in the Early Twentieth Century, London: UCL Press, 2018. Pp. 438. ISBN 978-1-78735-395-4. £50.00 (hardback). ISBN 978-1-78735-394-7. £30.00 (paperback). ISBN 978-1-78735-393-0 (open access). British Journal for the History of Science, 54(2):237–238, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/robert-bud-
paul-greenhalgh-frank-james-
and-morag-shiach-eds-being-
modern-the-cultural-impact-
of-science-in-the-early-
twentieth-century-london-
ucl-press-2018-pp-438-isbn-
9781787353954-5000-hardback-isbn-
9781787353947-3000-paperback-
isbn-9781787353930-open-access/
6F31B2333BB0A68D8D4A1910FDFF77D0.

Beiermann:2021:MST

[Bei21] Lea Beiermann. ‘A method for safe transmission’: the microscope slides of the American Postal Microscopical Club. British Journal for the History of Science, 54 (4):403–422, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
method-for-safe-
transmission-the-microscope-
slides-of-the-american-
postal-microscopical-club/
1269011CF843B0CC6758A38883ED6ABC.

Bekasova:2020:VTF

[Bek20] Alexandra Bekasova. Voyaging towards the future: the brig Rurik in the North Pacific and the emerging science of the sea. British Journal for the History of Science, 53 (4):469–495, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        voyaging-towards-the-future-
                                        the-brig-rurik-in-the-north-
                                        pacific-and-the-emerging-
                                        science-of-the-sea/
                                        52837ED83D7D9069F3B5B61465F0252D.

Bellis:2022:OSE

[Bel22] Richard T. Bellis. ‘The object of sense and experiment’: the ontology of sensation in William Hunter’s investigation of the human gravid uterus. British Journal for the History of Science, 55(2):227–246, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        object-of-sense-and-
                                        experiment-the-ontology-of-
                                        sensation-in-william-
                                        hunters-investigation-of-
                                        the-human-gravid-uterus/
                                        9C603438D58E747B15C1BFA50C2A533D.

Belteki:2023:WRC

[Bel23]  Daniel Belteki.  The winter

of raw computers: the history of the lunar and planetary reductions of the Royal Observatory, Greenwich. British Journal for the History of Science, 56(1):65–81, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.

cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
winter-of-raw-computers-the-
history-of-the-lunar-and-
planetary-reductions-of-the-
royal-observatory-greenwich/
A71791F9C2C0AB7125E49C7B56DFE5F1.

Bertozzi:2021:SOO

[Ber21] Eugenio Bertozzi. ‘Seeing with one’s own eyes’ and speaking to the mind: a history of the Wilson cloud chamber in the teaching of physics. British Journal for the History of Science, 54(2):177– 193, June 2021. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/seeing-with-
ones-own-eyes-and-speaking-
to-the-mind-a-history-of-
the-wilson-cloud-chamber-in-
the-teaching-of-physics/
8B43708F0415BFE3941805F7328D6ADB.

Bienias:2020:PEG

[Bie20] Barbara Bienias. The place of Edward Gresham’s Astrostereon (1603) in the discussion on cosmology and the Bible in the early modern period. British Journal for the History of Science, 53(4): 417–442, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
place-of-edward-greshams-

                                        astrostereon-1603-in-the-
                                        discussion-on-cosmology-and-
                                        the-bible-in-the-early-
                                        modern-period/
                                        4A0BBEA8D46280AC9C28FA0639A8BA2E.

Bigg:2023:CSM

[Big23] Charlotte Bigg. Communicating science, mediating presence: reflections on the present, past and future of conferencing. British Journal for the History of Science, 56(4):567–577, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        communicating-science-
                                        mediating-presence-
                                        reflections-on-the-present-
                                        past-and-future-of-
                                        conferencing/
                                        E747A8A7BE94485649E8466D279A6C58.

Bilgili:2020:RUH

[Bil20] Alper Bilgili. Retrospectives: Uses of history of science in the late Ottoman Empire and early republican Turkey. British Journal for the History of Science, 53 (1):109–117, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        retrospectives-uses-of-
                                        history-of-science-in-the-
                                        late-ottoman-empire-and-
                                        early-republican-turkey/
                                        FFDECA807EDF80ED84E638B221B19D54.

Bimm:2020:BRP

[Bim20] Jordan Bimm. Book review:

Philip W. Clements, Science in an Extreme Environment: The 1963 American Mount Everest Expedition. Pittsburgh: University of Pittsburgh Press, 2018. Pp. xvii + 269. ISBN 978-0-8229-4511-

6. $39.95 (paperback). British Journal for the History of Science, 53(1):121–123, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
philip-w-clements-science-
in-an-extreme-environment-
the-1963-american-mount-
everest-expedition-
pittsburgh-university-of-
pittsburgh-press-2018-pp-
xvii-269-isbn-9780822945116-
3995-paperback/756229A3BEE4019305B76C757EE1A605.

Blakley:2020:BRW

[Bla20] Christopher M. Blakley. Book review: Wendy Gonaver, The Peculiar Institution and the Making of Modern Psychiatry, 1840–1880. Chapel Hill: University of North Carolina Press, 2019. ISBN 978-1-4696-4844-6. $32.95 (paperback). British Journal for the History of Science, 53(4):596–597, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/wendy-gonaver-the-

                                        peculiar-institution-and-
                                        the-making-of-modern-
                                        psychiatry-18401880-chapel-
                                        hill-university-of-north-
                                        carolina-press-2019-isbn-
                                        9781469648446-3295-
                                        paperback/5AB051FF8ECEB727321707156A0715E6.

Bolman:2023:WML

[Bol23] Brad Bolman. What mysteries lay in spore: taxonomy, data, and the internationalization of mycology in Saccardo’s Sylloge Fungorum. British Journal for the History of Science, 56(3):369–390, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/what-mysteries-lay-
                                        in-spore-taxonomy-data-and-
                                        the-internationalization-of-
                                        mycology-in-saccardos-
                                        sylloge-fungorum/
                                        621C128689D7D3A31788AFEB40C39BF3.

Boon:2023:BRG

[Boo23a] Tim Boon. Book review: Grant Ramsey and Andreas De Block (eds), The Dynamics of Science: Computational Frontiers in History and Philosophy of Science. Pittsburgh: University of Pittsburgh Press, 2022. Pp. 308. ISBN 978-0-8229-4737-0. $60.00 (hardcover). British Journal for the History of Science, 56(4):584–585, December 2023. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-

journal-for-the-history-of-
science/article/grant-
ramsey-and-andreas-de-block-
eds-the-dynamics-of-science-
computational-frontiers-in-
history-and-philosophy-of-
science-pittsburgh-
university-of-pittsburgh-
press-2022-pp-308-isbn-9780822947370-
6000-hardcover/8331671B7B0B20EB4058E51DEB780209.

Boon:2023:PAS

[Boo23b] Tim Boon. Presidential address:

‘Some years of cudgelling my brains about the nature and function of science museums’: Frank Sherwood Taylor and the public role of the history of science. British Journal for the History of Science, 56(3):283–307, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/presidential-
address-some-years-of-
cudgelling-my-brains-about-
the-nature-and-function-of-
science-museums-frank-
sherwood-taylor-and-the-
public-role-of-the-history-
of-science/0904E615DFB0335C3DD1FA3104FF6D14.

Bowler:2022:BRL

[Bow22] Peter J. Bowler. Book reviews:

Lukas Rieppel, Assembling the Dinosaur: Fossil Hunters, Tycoons, and the Making of a Spectacle, Cambridge, MA: Harvard University Press, 2019. Pp. 325. ISBN 978-0-6747-3758-7. £23.95 (hardback). — Michael J. Benton, The Dinosaurs: How a Scientific Revolution Is Rewriting History, London: Thames and Hudson, 2020. Pp. 320. ISBN 978-0-5000-5200-6. £10.99 (paperback). British Journal for the History of Science, 55 (4):529–531, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
lukas-rieppel-assembling-
the-dinosaur-fossil-hunters-
tycoons-and-the-making-of-a-
spectacle-cambridge-ma-
harvard-university-press-
2019-pp-325-isbn-
9780674737587-2395-hardback-
michael-j-benton-the-dinosaurs-
how-a-scientific-revolution-is-
rewriting-history-london-thames-
and-hudson-2020-pp-320-isbn-9780500052006-
1099-paperback/95B884C81AF2EA08943DCF8F5114EAED.

Bowler:2023:BRC

[Bow23a] Peter J. Bowler. Book review: Claudine Cohen, Nos ancêtres dans les arbres: Penser l’évolution humaine. Paris: Editions du Seuil, 2021. Pp. 319. ISBN 978-2-0211-7599-8. EUR23.00 (paperback). British Journal for the History of Science, 56(1):128–129, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
claudine-cohen-nos-ancetres-
dans-les-arbres-penser-

levolution-humaine-paris-
editions-du-seuil-2021-pp-
319-isbn-9782021175998-2300-
paperback/
1B9E1DC71BD23A28E7D221B034F6FF6A.

Bowler:2023:BRF

[Bow23b] Peter J. Bowler. Book review: Frederick Burkhardt, James Secord and the editors of the Darwin Correspondence Project (eds.), The Correspondence of Charles Darwin, volume 22, 1874 Cambridge: Cambridge University Press, 2015. Pp. 855. ISBN 978-1-107-08872-6. £105.00 (hardback). British Journal for the History of Science, 56(3):409–410, September 2023. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/frederick-
burkhardt-james-secord-and-
the-editors-of-the-darwin-
correspondence-project-eds-
the-correspondence-of-
charles-darwin-volume-22-
1874-cambridge-cambridge-
university-press-2015-pp-855-
isbn-9781107088726-10500-hardback/
BAD36B9D7EA0A808041FCBFA6B0C26D8.

Bowler:2024:EE

[Bow24] Peter J. Bowler. The end of an era. British Journal for the History of Science, 57(1):113–117, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-

                                        for-the-history-of-science/
                                        article/end-of-an-era/
                                        9C5CAEA1ACC3225117C1C219F15F574E.

Bernhardt-Radu:2023:BRD

[BR23] Stefan Bernhardt-Radu. Book review: Daniel S. Milo, Good Enough: The Tolerance for Mediocrity in Nature and Society. Cambridge, MA: Harvard University Press, 2019. Pp. 310. ISBN 978-0-6745-0462-2. $28.95 (hardback). British Journal for the History of Science, 56(1):117–119, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/daniel-s-milo-good-
                                        enough-the-tolerance-for-
                                        mediocrity-in-nature-and-
                                        society-cambridge-ma-
                                        harvard-university-press-
                                        2019-pp-310-isbn-
                                        9780674504622-2895-hardback/
                                        273C7DDDED760004C87A561DF0ADB8E7.

Brock:2020:BRM

[Bro20] William H. Brock. Book review: Michael D. Barton, Janet Browne, Ken Corbett and Norman McMillan (eds.), The Correspondence of John Tyndall, vol. 6: The Correspondence, November 1856-February 1859. Pittsburgh, PA: University of Pittsburgh Press, 2018. Pp. lviii + 537. ISBN 978-0-8229-4533-8. $125.00. (hardback). British Journal for the History of Science, 53(4):598–599, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/michael-d-barton-
janet-browne-ken-corbett-
and-norman-mcmillan-eds-the-
correspondence-of-john-
tyndall-vol-6-the-
correspondence-november-
1856february-1859-
pittsburgh-pa-university-of-pittsburgh-
press-2018-pp-lviii-537-isbn-
9780822945338-12500-hardback/
6926EABB3ABCCF08F8278192EB75CFF4.

Bigg:2023:AGH

[BRSW23] Charlotte Bigg, Jessica Reinisch, Geert Somsen, and Sven Widmalm. The art of gathering: histories of international scientific conferences. British Journal for the History of Science, 56(4): 423–433, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
art-of-gathering-histories-
of-international-scientific-
conferences/
69855DA669EB68A587CCD85977C07878.

Bruinsma:2022:BRV

[Bru22] Mette Bruinsma. Book review:

Veronica Della Dora, The Mantle of the Earth: Genealogies of a Geographical Metaphor. Chicago: The University of Chicago Press, 2021. Pp. 416. ISBN 978-0-226-74129-1. $65.00 (hardback). British Journal for the History of Science, 55 (2):249–250, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/veronica-
                                        della-dora-the-mantle-of-
                                        the-earth-genealogies-of-a-
                                        geographical-metaphor-
                                        chicago-the-university-of-
                                        chicago-press-2021-pp-416-
                                        isbn-9780226741291-6500-hardback/
                                        C896AA3CB8B66787A260726FD3D266FC.

Bruyninckx:2024:BRF

[Bru24] Joeri Bruyninckx. Book review: Fanny Gribenski, Tuning the World: The Rise of 440 Hertz in Music, Science, and Politics, 1859–1955, Chicago: University of Chicago Press, 2023. Pp. 280. ISBN 978-0-226-82326-3. $55.00 (cloth). British Journal for the History of Science, 57(1):127–129, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/fanny-gribenski-
                                        tuning-the-world-the-rise-
                                        of-440-hertz-in-music-
                                        science-and-politics-
                                        18591955-chicago-university-
                                        of-chicago-press-2023-pp-
                                        280-isbn-9780226823263-5500-
                                        cloth/0A00F4BE95C4FB4CF48EEDB5FB9C9D4B.

Boon:2020:TBO

[BS20] Tim Boon and Charlotte Sleigh. Two BSHS online alternatives to conventional conferences. British Journal for the History of Science, 53(4):553–554, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
two-bshs-online-
alternatives-to-
conventional-conferences/
8149048F6884315513E5231A47EEC67A.

Bucciarelli:2021:BRJ

[Buc21] Lucia Bucciarelli. Book review:

J. B. Shank, Before Voltaire: The French Origins of ‘Newtonian’ Mechanics, 1680–1715. Chicago: The University of Chicago Press, 2018. Pp. 464. ISBN: 978-0-226-50929-7. $55.00 (paperback). British Journal for the History of Science, 54 (3):396–398, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/jb-shank-
before-voltaire-the-french-
origins-of-newtonian-
mechanics-16801715-chicago-
the-university-of-chicago-
press-2018-pp-464-isbn-
9780226509297-5500-
paperback/6146D818DCB238492D873694768018F8.

Bud:2024:BRM

[Bud24] Robert Bud. Book review: Maria Rentetzi, Seduced by Radium: How Industry Transformed Science in the American Marketplace, Pittsburgh, PA: University of Pittsburgh Press, 2022. Pp. xi + 292. ISBN 978-0-8229-4706-6. $35.00 (hardback). British Journal for the History of Science, 57 (1):119–121, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/maria-
                                        rentetzi-seduced-by-radium-
                                        how-industry-transformed-
                                        science-in-the-american-
                                        marketplace-pittsburgh-pa-
                                        university-of-pittsburgh-
                                        press-2022-pp-xi-292-isbn-
                                        9780822947066-3500-hardback/241891E696E363ACC1129E72206662B6.

Burton:2022:CGB

[Bur22] Elise K. Burton. Comparative globalizations: building and dismantling genetic laboratories in Lebanon. British Journal for the History of Science, 55(4):495–513, December 2022. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/comparative-
                                        globalizations-building-and-
                                        dismantling-genetic-
                                        laboratories-in-lebanon/
                                        5E248827702CC0CC79D148EB7C7E61FF.

Buschmann:2022:CDD

[Bus22] Rainer F. Buschmann. Contested duplicates: disputed negotiations surrounding ethnographic doppelgängers in German New Guinea, 1898–1914. British Journal for the History of Science, 55(3):297–318, September 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
contested-duplicates-
disputed-negotiations-
surrounding-ethnographic-
doppelgangers-in-german-new-
guinea-18981914/
BCEAE036B3140E806AA6D03309CAA154.

Carson:2021:AMN

[Car21] Sarah Carson. Anticipating the monsoon: the necessity and impossibility of the seasonal weather forecast for South Asia, 1886– 1953. British Journal for the History of Science, 54(3):305–325, September 2021. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
anticipating-the-monsoon-
the-necessity-and-
impossibility-of-the-
seasonal-weather-forecast-
for-south-asia-18861953/
1FC69001769FA9B58A1CFA7A69E167A2.

Chu:2023:BRB

[Chu23] Leo Chu. Book review: Ben Nobbs-Thiessen, Landscape of Migration: Mobility and Environmental Change on Bolivia’s Tropical Frontier, 1952 to the Present. Chapel Hill, NC: University of North Carolina Press, 2020. Pp. 342. ISBN 978-1-4696-5609-0. $99.00 (hardback). British Journal for the History of Science, 56(1):124–126, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
ben-nobbsthiessen-landscape-
of-migration-mobility-and-
environmental-change-on-
bolivias-tropical-frontier-
1952-to-the-present-chapel-
hill-nc-university-of-north-
carolina-press-2020-pp-342-isbn-
9781469656090-9900-hardback/E7278E0324BA70E0631B3FB2D8D48ECC.

Cozzoli:2020:BRM

[Coz20] Daniele Cozzoli. Book review: Marcos Cueto, Theodore M. Brown and Elizabeth Fee, The World Health Organization: A History. Cambridge: Cambridge University Press, 2019. Pp. ix + 373. ISBN 978-1-108-72884-

3. £26.99 (paperback). British Journal for the History of Science, 53(1):126–128, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
marcos-cueto-theodore-m-
brown-and-elizabeth-fee-the-
world-health-organization-a-
history-cambridge-cambridge-
university-press-2019-pp-ix-
373-isbn-9781108728843-2699-
paperback/BE3BBB107CCAA90F39AE8C7A84CDB8D8.

deAsua:2020:SCP

[dA20] Miguel de Asúa. Science, Catholicism and politics in Argentina (1910–1935). British Journal for the History of Science, 53(2):139– 158, June 2020. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/science-
catholicism-and-politics-in-
argentina-19101935/
31F7C79840A424B3E2D65A9BA0DD7713.

Davies:2021:PMR

[Dav21] Stephen M. Davies. Priorities in medical research: elite dynamics in a pivotal episode for British health research. British Journal for the History of Science, 54(2):195–211, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
priorities-in-medical-
research-elite-dynamics-in-
a-pivotal-episode-for-
british-health-research/
305F78561AB890CB7BD5CFCAF5F2E751.

deBont:2020:EGP

[dB20] Raf de Bont. Eating game: proteins, international conservation and the rebranding of African wildlife, 1955–1965. British Journal for the History of Science, 53(2):183–205, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
eating-game-proteins-
international-conservation-
and-the-rebranding-of-
african-wildlife-19551965/
50786E80B0037B1BBEDF9250988DB00C.

Deen:2020:BRL

[Dee20] Rosa Deen. Book review: Lorei Gruen (ed.), Critical Terms for Animal Studies. Chicago: The University of Chicago Press, 2018. Pp. 472. ISBN 978-0-226-35542-

9. £32.50/£24.50 (paperback). British Journal for the History of Science, 53(1):132–134, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
lorei-gruen-ed-critical-
terms-for-animal-studies-
chicago-the-university-of-
chicago-press-2018-pp-472-
isbn-9780226355429-32502450-
paperback/
68713C71201FDEE538C4B2C00293F23F.

Dessain:2020:BRI

[Des20] H.-F. Dessain. Book review: Ian Duncan, Human Forms: The Novel in the Age of Evolution. Princeton, NJ: Princeton University Press, 2019. Pp. 304. ISBN 978-0-691-17507-2. £30.00 (hardcover). British Journal for the History of Science, 53(2):285–287, June 2020. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/ian-duncan-human-
forms-the-novel-in-the-age-
of-evolution-princeton-nj-
princeton-university-press-
2019-pp-304-isbn-
9780691175072-3000-
hardcover/
9C9BBB377412CE54508746DE4BDA1AE3.

Dessain:2021:BRT

[Des21] H.-F. Dessain. Book review:

Tom Williamson, Humphry Repton: Landscape Design in an Age of Revolution London: Reaktion Books, 2020. Pp. 312. ISBN 978-1-78914-299-0. £35.00 (hardcover). British Journal for the History of Science, 54(1):115–116, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
tom-williamson-humphry-
repton-landscape-design-in-
an-age-of-revolution-london-
reaktion-books-2020-pp-312-
isbn-9781789142990-3500-
hardcover/
1851102F7BB8358AD72022A2A49A7B7A.

DiCastri:2023:BRP

[Di 23] Theo Di Castri. Book reviews:

Philippe Fontaine and Jefferson D. Pooley, Society on the Edge: Social Science and Public Policy in the Postwar United States. Cambridge: Cambridge University Press, 2020. Pp. 280. ISBN 978-1-1084-8713-9. £74.99 (hardback). ISBN 978-1-1087-3219-2. £26.99 (paperback). British Journal for the History of Science, 56(1):132– 134, March 2023. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/philippe-
fontaine-and-jefferson-d-
pooley-society-on-the-edge-
social-science-and-public-
policy-in-the-postwar-
united-states-cambridge-
cambridge-university-press-
2020-pp-280-isbn-9781108487139-
7499-hardback-isbn-9781108732192-
2699-paperback/40F69300EB1735D15A27A21AFF2D3378.

Dishington:2024:VTM

[Dis24] Rachel Dishington. ‘The very term mensuration sounds engineer-like’: measurement and engineering authority in nineteenth-century river management. British Journal for the History of Science, 57(1):21– 41, March 2024. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/very-term-
mensuration-sounds-
engineerlike-measurement-
and-engineering-authority-
in-nineteenthcentury-river-
management/
3E96E060014B10F4D73B61234406F961.

Doley:2021:BRT

[Dol21] Caitlin Doley. Book review: Tom Stammers, The Purchase of the Past: Collecting Culture in Postrevolutionary Paris c. 1790–1890. Cambridge: Cambridge University Press, 2020. Pp. 370. ISBN: 978-1-108-78126-8. £90.00 (hardback). British Journal for the History of Science, 54(3):398–399, September 2021. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/tom-
stammers-the-purchase-of-
the-past-collecting-culture-
in-postrevolutionary-paris-
c17901890-cambridge-
cambridge-university-press-
2020-pp-370-isbn-
9781108781268-9000-hardback/E6C3C9352EF1C5D65EE8F1867BF38813.

Duancic:2021:BRA

[Dua21] Vedran Duancić. Book reviews:

Audra Wolfe, Freedom’s Laboratory: The Cold War Struggle for the Soul of Science. Baltimore: Johns Hopkins University Press, 2018. Pp. x + 302. ISBN 978-1-4214-2673-0. $29.95 (hardback). ISBN 978-1-4214-3908-2. $19.95 (paperback). British Journal for the History of Science, 54(1):120– 122, March 2021. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/audra-wolfe-

                                        freedoms-laboratory-the-
                                        cold-war-struggle-for-the-
                                        soul-of-science-baltimore-
                                        johns-hopkins-university-
                                        press-2018-pp-x-302-isbn-
                                        9781421426730-2995-hardback-
                                        isbn-9781421439082-1995-paperback/
                                        F3E75245BC3CB663E7B5314B3AC1FFAF.

Dunn:2022:BRJ

[Dun22] Richard Dunn. Book review: Jennifer M. Rampling, The Experimental Fire: Inventing English Alchemy, 1300– 1700, Chicago and London: The University of Chicago Press, 2020. Pp. 416. ISBN 978-0-226-71070-9. £28.00/$35.00 (hardback). British Journal for the History of Science, 55(1):119–120, March 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        jennifer-m-rampling-the-
                                        experimental-fire-inventing-
                                        english-alchemy-13001700-
                                        chicago-and-london-the-
                                        university-of-chicago-press-
                                        2020-pp-416-isbn-
                                        9780226710709-28003500-hardback/
                                        4E90AE5EF256E7C9949EAF260A8CAA82.

Dunning:2023:BRA

[Dun23] David E. Dunning. Book review: Andrew Fiss, A History of Communication and Anxiety in the American Mathematics Classroom, New Brunswick, NJ: Rutgers University Press, 2020. Pp. 216. ISBN 978-1-9788-2021-0. $150.00 (cloth). British Journal for the History of Science, 56(3): 411–412, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/andrew-fiss-
a-history-of-communication-
and-anxiety-in-the-american-
mathematics-classroom-new-
brunswick-nj-rutgers-
university-press-2020-pp-
216-isbn-9781978820210-
15000-cloth/C5321F2F19886C777235023C0FA6A49B.

Durlo:2023:BRK

[Dur23] Andrea Durlo. Book review:

Kevin McCain and Kostas Kampourakis: What Is Scientific Knowledge? An Introduction to Contemporary Epistemology of Science. London: Routledge, 2019. Pp. 328. ISBN 978-1-1385-7015-

3. £36.99 (paperback). British Journal for the History of Science, 56(1):131–132, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
kevin-mccain-and-kostas-
kampourakis-what-is-
scientific-knowledge-an-
introduction-to-
contemporary-epistemology-
of-science-london-routledge-
2019-pp-328-isbn-9781138570153-
3699-paperback/267DEB1282AB77D7B98B0B344CB16DAA.

Diogo:2023:CDV

[DUS23] Maria Paula Diogo, Paula Urze, and Ana Simões. Cartoon diplomacy: visual strategies, imperial rivalries and the 1890 British Ultimatum to Portugal. British Journal for the History of Science, 56:147–166, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/cartoon-
                                        diplomacy-visual-strategies-
                                        imperial-rivalries-and-the-
                                        1890-british-ultimatum-to-
                                        portugal/
                                        40313FFA1FE96BC0EEC321285D61D02A.

Eichberg:2021:BRV

[Eic21] Stephanie Eichberg. Book review: Vaughn Scribner, Merpeople: A Human History. London: Reaktion Books, 2020. Pp. 320. ISBN: 978-1-78914-314-0. £20.00 (hardback). British Journal for the History of Science, 54(1):122–124, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/vaughn-scribner-
                                        merpeople-a-human-history-
                                        london-reaktion-books-2020-
                                        pp-320-isbn-9781789143140-
                                        2000-hardback/
                                        4A9AFBA084576FC81CED7EBB49EE03F7.

Eriksen:2023:CFB

[EW23] Christoffer Basse Eriksen and Xinyi Wen. Colouring flowers: books, art, and experiment in the household of Margery and Henry Power. British Journal for the History of Science, 56(1):21–43, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/colouring-flowers-
books-art-and-experiment-in-
the-household-of-margery-
and-henry-power/
6BA23F60CC9962D8045AF81F8B75799D.

Exley:2023:BRC

[Exl23] Grace Exley. Book review: Claire G. Jones, Alison E. Martin and Alexis Wolf (eds), The Palgrave Handbook of Women and Science since 1660. London: Palgrave Macmillan, 2022. Pp. 658. ISBN 978-3-030-78972-5. £149.99 (hardcover). British Journal for the History of Science, 56(4): 586–587, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
claire-g-jones-alison-e-
martin-and-alexis-wolf-eds-
the-palgrave-handbook-of-
women-and-science-since-
1660-london-palgrave-
macmillan-2022-pp-658-isbn-
9783030789725-14999-hardcover/
19EFB0C2DD6B80D23E3049786F8D349F.

Fara:2020:BRB

[Far20] Patricia Fara. Book review: Ben Marsh, Unravelled Dreams: Silk and the Atlantic World, 1500– 1840. Cambridge: Cambridge University Press, 2020. Pp. xiv + 487. ISBN 978-1-108-41828-7. £29.99 (hardback). British Journal for the History of Science, 53(4): 594–596, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
ben-marsh-unravelled-dreams-
silk-and-the-atlantic-world-
15001840-cambridge-
cambridge-university-press-
2020-pp-xiv-487-isbn-
9781108418287-2999-hardback/
DBA61DF3CF2542AA430033C8AF332325.

Fara:2021:BRZ

[Far21] Patricia Fara. Book review: Zachary Dorner, Merchants of Medicines: The Commerce and Coercion of Health in Britain’s Long Eighteenth Century. Chicago: The University of Chicago Press, 2020. Pp. 280. ISBN: 978-0-226-70680-1. $50.00 (hardback). British Journal for the History of Science, 54 (4):531–532, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
zachary-dorner-merchants-of-
medicines-the-commerce-and-
coercion-of-health-in-
britains-long-eighteenth-
century-chicago-the-
university-of-chicago-press-

2020-pp-280-isbn-9780226706801-
5000-hardback/2F5D8423E34CC446A0941566CA1457BD.

Fara:2023:BRJ

[Far23]  Patricia Fara.  Book review:

Jacqueline Mitton and Simon Mitton, Vera Rubin: A Life. Cambridge, MA: Belknap Press, 2021. Pp. x + 309. ISBN 978-0-674-91919-8. £23.95 (hardback). British Journal for the History of Science, 56:278–279, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
jacqueline-mitton-and-simon-
mitton-vera-rubin-a-life-
cambridge-ma-belknap-press-
2021-pp-x-309-isbn-
9780674919198-2395-hardback/
C001FF83EB9177B6468D91F6062CDDB2.

Figueroa:2021:BRA

[Fig21] Javier Poveda Figueroa. Book reviews: Alexander Wragge-Morley, Aesthetic Science: Representing Nature in the Royal Society of London, 1650–1720. Chicago: The University of Chicago Press, 2020. Pp. 272. ISBN: 978-0-226-68086-6. $120.00 (hardback). ISBN: 978-0-226-68072-9. $40.00 (paperback). British Journal for the History of Science, 54(3):394–396, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/alexander-

                                        wraggemorley-aesthetic-
                                        science-representing-nature-
                                        in-the-royal-society-of-
                                        london-16501720-chicago-the-
                                        university-of-chicago-press-
                                        2020-pp-272-isbn-
                                        9780226680866-12000-hardback-
                                        isbn-9780226680729-4000-paperback/
                                        3CA2611693D97FAF422E9F0BEC55AF6B.

Forster:2023:SSC

[For23] Laura C. Forster. Sex, science and curated community at the World League for Sexual Reform 1929 conference. British Journal for the History of Science, 56(4):469– 484, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/sex-science-
                                        and-curated-community-at-
                                        the-world-league-for-sexual-
                                        reform-1929-conference/
                                        B08C19B5A69BBDCAC1A42D0ECC414B3F.

Frampton:2020:BRJ

[Fra20] Sally Frampton. Book review: Jacqueline H. Wolf, Cesarean Section: An American History of Risk, Technology, and Consequence. Baltimore: Johns Hopkins University Press, 2018. Pp. 320. ISBN 978-1-4214-2552-8. $49.95 (hardback). British Journal for the History of Science, 53(1):136– 137, March 2020. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-

science/article/jacqueline-
h-wolf-cesarean-section-an-
american-history-of-risk-
technology-and-consequence-
baltimore-johns-hopkins-
university-press-2018-pp-
320-isbn-9781421425528-4995-
hardback/B390AD3BFDC91541E7201878B7D1E65B.

Frank:2021:FAO

[Fra21] Mark E. Frank. Frontier atmosphere: observation and regret at Chinese weather stations in Tibet, 1939–1949. British Journal for the History of Science, 54(3): 361–379, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/frontier-
atmosphere-observation-and-
regret-at-chinese-weather-
stations-in-tibet-19391949/
355E8143AB3B26F40222D703084568EA.

Grevsmuhl:2023:SIT

[GB23] Sebastian V. Grevsmühl and Régis Briday. Satellite images as tools of visual diplomacy: NASA’s ozone hole visualizations and the Montreal Protocol negotiations. British Journal for the History of Science, 56:247–267, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
satellite-images-as-tools-
of-visual-diplomacy-nasas-
ozone-hole-visualizations-

                                        and-the-montreal-protocol-
                                        negotiations/
                                        EA86D11886E97CBABCDFA42126184129.

Gillin:2020:TTR

[Gil20] Edward J. Gillin. Tremoring transits: railways, the Royal Observatory and the capitalist challenge to Victorian astronomical science. British Journal for the History of Science, 53(1):1–24, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        tremoring-transits-railways-
                                        the-royal-observatory-and-
                                        the-capitalist-challenge-to-
                                        victorian-astronomical-
                                        science/
                                        3EA125839341F1C175D83538F3D9BA78.

Goodrum:2020:BRE

[Goo20] Matthew R. Goodrum. Book review: Erika Lorraine Milam, Creatures of Cain: The Hunt for Human Nature in Cold War America. Princeton, NJ: Princeton University Press, 2019. Pp. 408. ISBN 978-0-691-18188-2. $29.95 (hardcover). British Journal for the History of Science, 53(4):591–593, December 2020. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/erika-
                                        lorraine-milam-creatures-of-
                                        cain-the-hunt-for-human-
                                        nature-in-cold-war-america-

princeton-nj-princeton-
university-press-2019-pp-
408-isbn-9780691181882-2995-
hardcover/736A38931501C3A93F0798782C96F6EE.

Gordin:2022:ELM

[Gor22] Michael D. Gordin. Einsteinian language: Max Talmey, Benjamin Lee Whorf and linguistic relativity. British Journal for the History of Science, 55(2):145–165, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
einsteinian-language-max-
talmey-benjamin-lee-whorf-
and-linguistic-relativity/
ECE744AA2067BA1EC96296C2E89DAC18.

Grodwohl:2023:BRM

[Gro23] Jean-Baptiste Grodwohl. Book review: Michel Morange, The Black Box of Biology: A History of the Molecular Revolution. Cambridge, MA: Harvard University Press, 2020. Pp. 528. ISBN 978-0-674-28136-3. £40.95 (hardcover). British Journal for the History of Science, 56(4): 588–589, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
michel-morange-the-black-
box-of-biology-a-history-of-
the-molecular-revolution-
cambridge-ma-harvard-
university-press-2020-pp-

                                        528-isbn-9780674281363-4095-
                                        hardcover/B2A4C3723C18910BF9BBE1DB6E245F2C.

Gross:2024:BRC

[Gro24] Benjamin Gross. Book review: Cyrus C. M. Mody, The Squares: US Physical and Engineering Scientists in the Long 1970s, Cambridge, MA: MIT Press, 2023. Pp. 422. ISBN 978-0-262-54361-

3. $65.00 (paperback). British Journal for the History of Science, 57(1):121–123, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        cyrus-cm-mody-the-squares-
                                        us-physical-and-engineering-
                                        scientists-in-the-long-
                                        1970s-cambridge-ma-mit-
                                        press-2023-pp-422-isbn-
                                        9780262543613-6500-
                                        paperback/
                                        0CF21C826A2FB4D2C4830B783A306DF0.

Giovannetti-Singh:2021:BRP

[GS21] Gianamar Giovannetti-Singh. Book review: Pratik Chakrabarti, Inscriptions of Nature: Geology and the Naturalization of Antiquity. Baltimore: Johns Hopkins University Press, 2020. Pp. 280. ISBN: 987-1-4214-3874-0. $54.95 (hardback). British Journal for the History of Science, 54(1):118–120, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-

science/article/pratik-
chakrabarti-inscriptions-of-
nature-geology-and-the-
naturalization-of-antiquity-
baltimore-johns-hopkins-
university-press-2020-pp-
280-isbn-9817421438740-5495-hardback/
DA7541550E65A554C49E11DA3EEFD8A9.

The title and URL have an incorrect ISBN-13 prefix: 987 and 9817 should both be 978. The correct ISBN-13 is 978-1-4214-3874-0 (hardcover).

Giovannetti-Singh:2022:BRW

[GS22]  Gianamar  Giovannetti-Singh.

Book review: William Beinart and Saul Dubow, The Scientific Imagination in South Africa: 1700 to the Present, Cambridge: Cambridge University Press, 2022, pp. 406. ISBN 987-1-1088-3708-8. £64.99 (hardback). British Journal for the History of Science, 55 (1):121–122, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/william-
beinart-and-saul-dubow-the-
scientific-imagination-in-
south-africa-1700-to-the-
present-cambridge-cambridge-
university-press-2022-pp-
406-isbn-9871108837088-6499-hardback/
19A29A988363D007DAFB5B894027D25F.

The title and URL have an invalid ISBN-13 prefix: 987 should be 978. The correct ISBN-13 values are 978-1-108-83708-8 (hardcover) and 978-1-108-94008-5 (paperback).

Guffroy:2021:BRM

[Guf21] Yohann Guffroy. Book reviews: Malcolm Dick and Caroline Archer– Parré (eds.), James Watt, 1736– 1819: Culture, Innovation, Enlightenment. Liverpool: Liverpool University Press, 2020. Pp. 280. ISBN: 978-1-78962-081-8. £80.00 (hardback). ISBN: 978-1-78962-082-5. £24.99 (paperback). British Journal for the History of Science, 54(3):400–401, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        malcolm-dick-and-caroline-
                                        archerparre-eds-james-watt-
                                        17361819-culture-innovation-
                                        enlightenment-liverpool-
                                        liverpool-university-press-
                                        2020-pp-280-isbn-
                                        9781789620818-8000-hardback-isbn-
                                        9781789620825-2499-paperback/
                                        8A972503C086D2856849735E58A5E44D.

Honeybun-Arnolda:2020:BRM

[HA20] Elliot Honeybun-Arnolda. Book review: Michael Oppenheimer, Naomi Oreskes, Dale Jamieson, Keynyn Brysse, Jessica O’reilly, Matthew Shindell and Milena Wazeck, Discerning Experts: The Practices of Scientific Assessment for Environmental Policy. Chicago and London: The University of Chicago Press, 2019. Pp. ix + 281. ISBN 978-0-226-60201-1. $35.00 (paperback). British Journal for the History of Science, 53(1):128– 129, March 2020. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/michael-
oppenheimer-naomi-oreskes-
dale-jamieson-keynyn-brysse-
jessica-oreilly-matthew-
shindell-and-milena-wazeck-
discerning-experts-the-
practices-of-scientific-
assessment-for-environmental-
policy-chicago-and-london-the-
university-of-chicago-press-2019-
pp-ix-281-isbn-9780226602011-
3500-paperback/C0FFD7F2846745895EE20848B8F81957.

Honeybun-Arnolda:2021:BRE

[HA21] Elliott Honeybun-Arnolda. Book reviews: Etienne S. Benson, Surroundings: A History of Environments and Environmentalisms. Chicago: The University of Chicago Press, 2020. Pp. 296. ISBN: 978-0-226-70615-3. $82.50 (hardback). ISBN: 978-0-226-70629-0. $27.50 (paperback). British Journal for the History of Science, 54(3):392–394, September 2021. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/etienne-s-
benson-surroundings-a-
history-of-environments-and-
environmentalisms-chicago-
the-university-of-chicago-

                                        press-2020-pp-296-isbn-
                                        9780226706153-8250-hardback-isbn-
                                        9780226706290-2750-paperback/
                                        C33FA7C6C8E158939C3CC68E7F711562.

Honeybun-Arnolda:2024:SES

[HA24] Elliot Honeybun-Arnolda. Scientizing the ‘environment’: Solly Zuckerman and the idea of the School of Environmental Sciences. British Journal for the History of Science, 57(1):99–112, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        scientizing-the-environment-
                                        solly-zuckerman-and-the-
                                        idea-of-the-school-of-
                                        environmental-sciences/
                                        9A0307D4D7E158EC7332F6482B11443C.

Haberer:2024:BRV

[Hab24] Maximilian Haberer. Book review: Victoria Tkaczyk, Thinking with Sound: A New Program in the Sciences and Humanities around 1900, Chicago: The University of Chicago Press, 2023. Pp. 304. ISBN 978-0-226-82328-7. $55.00 (cloth). British Journal for the History of Science, 57(1):129–131, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        victoria-tkaczyk-thinking-
                                        with-sound-a-new-program-in-
                                        the-sciences-and-humanities-

around-1900-chicago-the-
university-of-chicago-press-
2023-pp-304-isbn-
9780226823287-5500-cloth/
380C6DC3D9988C6DFE2E15E0D053B4B5.

Hanley:2020:BRA

[Han20] Anne Hanley. Book review:

Amelia Bonea, Melissa Dickson, Sally Shuttleworth and Jennifer Wallis, Anxious Times: Medicine and Modernity in Nineteenth-Century Britain. Pittsburgh: University of Pittsburgh Press, 2019. Pp. 320. ISBN 978-0-8229-4551-2, $50.00 (hardcover). British Journal for the History of Science, 53 (1):119–121, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/amelia-
bonea-melissa-dickson-sally-
shuttleworth-and-jennifer-
wallis-anxious-times-
medicine-and-modernity-in-
nineteenthcentury-britain-
pittsburgh-university-of-pittsburgh-
press-2019-pp-320-isbn-9780822945512-
5000-hardcover/DA18E80463DFEAE7F3271139BED89B19.

Hardy:2021:BRA

[Har21] Penelope K. Hardy. Book review: Antony Adler, Neptune’s Laboratory: Fantasy, Fear, and Science at Sea Cambridge, MA and London: Harvard University Press, 2019. Pp. 256. ISBN: 978-0-674-97201-8. £31.95 (hardback). British Journal for the History of Science, 54(1):111–113, March 2021. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/antony-adler-
                                        neptunes-laboratory-fantasy-
                                        fear-and-science-at-sea-
                                        cambridge-ma-and-london-
                                        harvard-university-press-
                                        2019-pp-256-isbn-
                                        9780674972018-3195-hardback/
                                        9717B6A0406F13136698DA170329DEC0.

Hayes:2021:AHC

[Hay21a] Emily Hayes. The Anthropocene & the humanities: From climate change to a new age of sustainability. British Journal for the History of Science, 54(3): 381–385, September 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        anthropocene-the-humanities-
                                        from-climate-change-to-a-
                                        new-age-of-sustainability/
                                        DC0FF14BFC0FE24FAD82B2EA5F26D7EF.

Hayes:2021:BRA

[Hay21b] Emily Hayes. Book review: Andrei Pop, A Forest of Symbols: Art, Science, and Truth in the Long Nineteenth Century. New York: Zone Books, 2019. Pp. 320. ISBN 978-1-935408-36-9. £25.00 (hardback). British Journal for the History of Science, 54(4): 536–539, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
andrei-pop-a-forest-of-
symbols-art-science-and-
truth-in-the-long-
nineteenth-century-new-york-
zone-books-2019-pp-320-isbn-
9781935408369-2500-hardback/
46871408D8490351C54A49F68FCA6973.

Hayes:2023:BRM

[Hay23] Emily Hayes. Book review:

Matthew H. Edney, Cartography: The Ideal and Its History. Chicago: The University of Chicago Press, 2019. Pp. 296. ISBN: 978-0-226-60568-5. $32.00 (paperback). British Journal for the History of Science, 56:269– 271, June 2023. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
matthew-h-edney-cartography-
the-ideal-and-its-history-
chicago-the-university-of-
chicago-press-2019-pp-296-
isbn-9780226605685-3200-
paperback/
0DBD2E8881ED082587CD0A3ABEBF065D.

Heering:2023:TMR

[Hee23] Peter Heering. Transformations:

the material representation of historical experiments in science teaching. British Journal for the History of Science, 56(3): 351–368, September 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        transformations-the-
                                        material-representation-of-
                                        historical-experiments-in-
                                        science-teaching/
                                        30408D9327C495A54C1DBBC1E390BF89.

Helsing:2020:BRM

[Hel20] Daniel Helsing. Book review: Michael Bravo, North Pole: Nature and Culture. London: Reaktion Books, 2019. Pp. 254. ISBN 978-1-78914-008-8. $24.95/£14.96 (paperback). British Journal for the History of Science, 53 (1):131–132, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        michael-bravo-north-pole-
                                        nature-and-culture-london-
                                        reaktion-books-2019-pp-254-
                                        isbn-9781789140088-24951496-
                                        paperback/
                                        37B3884E330C086F5FE7F651218FD375.

Hemmens:2022:BRM

[Hem22] Sue Hemmens. Book review: Michel Anctil, Luminous Creatures: The History and Science of Light Production in Living Organisms, Montreal and Kingston: McGill-Queen’s University Press, 2018. Pp. xvii + 467. ISBN: 978-0-7735-5312-5. $49.95 (hardback). British Journal for the History of Science, 55(4):527–529, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/michel-anctil-
luminous-creatures-the-
history-and-science-of-
light-production-in-living-
organisms-montreal-and-
kingston-mcgillqueens-
university-press-2018-pp-
xvii-467-isbn-9780773553125-4995-
hardback/2E3FE162F6076EE73D8AFE43E299D42D.

Heslop:2023:BRH

[Hes23] Beck Chamberlain Heslop. Book review: Hans Radder, From Commodification to the Common Good: Reconstructing Science, Technology and Society. Pittsburgh: University of Pittsburgh Press, 2019. Pp. 312. ISBN 978-0-8229-4579-6. $50.00 (hardcover). British Journal for the History of Science, 56:271–273, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/hans-radder-
from-commodification-to-the-
common-good-reconstructing-
science-technology-and-
society-pittsburgh-
university-of-pittsburgh-
press-2019-pp-312-isbn-
9780822945796-5000-hardcover/
EF17907F39AD6E900F73634610187881.

Houghton:2020:RSU

[HLDS20] Jemma Houghton, Alexander Longworth-Dunbar, and Nicola Sugden. ‘Research sharing’ using social media: online conferencing and the experience of #BSHSGlobalHist. British Journal for the History of Science, 53(4):555–573, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/research-sharing-
                                        using-social-media-online-
                                        conferencing-and-the-
                                        experience-of-
                                        bshsglobalhist/
                                        AB6B1C7AB14DA2EE2C827EC8EFD0D0BC.

Heumann:2022:IID

[HMB22] Ina Heumann, Anne Greenwood MacKinney, and Rainer Buschmann. Introduction: the issue of duplicates. British Journal for the History of Science, 55(3):257–278, September 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/introduction-the-
                                        issue-of-duplicates/
                                        B8811D4F80ED5CB40E357A41D4BBE5C2.

Holmes:2020:BRN

[Hol20] Tarquin Holmes. Book review: Nicole C. Nelson, Model Behavior: Animal Experiments, Complexity, and the Genetics of Psychiatric Disorders. Chicago and London: The University of Chicago Press, 2018. Pp. 255. ISBN 978-0-226-54608-7. $30.00 (paperback). British Journal for the History of Science, 53(2):279–280, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
nicole-c-nelson-model-
behavior-animal-experiments-
complexity-and-the-genetics-
of-psychiatric-disorders-
chicago-and-london-the-
university-of-chicago-press-
2018-pp-255-isbn-
9780226546087-3000-paperback/
64925BD46110A47C46DBD4689BF5BE4A.

Holmes:2021:BRV

[Hol21]  Tarquin Holmes.  Book re-

view: Vanessa Heggie, Higher and Colder: A History of Extreme Physiology and Exploration Chicago and London: The University of Chicago Press, 2019. Pp. 253. ISBN 978-0-226-65088-

3. $40.00 (hardback). British Journal for the History of Science, 54(2):239–240, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
vanessa-heggie-higher-and-
colder-a-history-of-extreme-
physiology-and-exploration-
chicago-and-london-the-
university-of-chicago-press-
2019-pp-253-isbn-
9780226650883-4000-hardback/DBD5292B2F66BA87E8413C4E3FADC455.

Holmes:2022:BRC

[Hol22]  Charlotte Holmes.  Book re-

view: Claire Hickman, The Doctor’s Garden: Medicine, Science, and Horticulture in Britain. New Haven: Yale University Press, 2021. Pp. xiv + 238. ISBN 978-0-300-23610-1. £30.00 (hardback). British Journal for the History of Science, 55(2):253–255, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        claire-hickman-the-doctors-
                                        garden-medicine-science-and-
                                        horticulture-in-britain-new-
                                        haven-yale-university-press-
                                        2021-pp-xiv-238-isbn-
                                        9780300236101-3000-hardback/
                                        F20C5EC56ECBE1F443E4003CB791D726.

Holland:2023:BRM

[Hol23a] Sarah Holland. Book review: Mikkel Borch-Jacobsen, Freud’s Patients: A Book of Lives. London: Reaktion Books, 2021. Pp. 304. ISBN 978-1-7891-4455-0. £20.00 (hardback). British Journal for the History of Science, 56(3):413–414, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        mikkel-borchjacobsen-freuds-
                                        patients-a-book-of-lives-
                                        london-reaktion-books-2021-
                                        pp-304-isbn-9781789144550-

2000-hardback/
011EE83E2A71B945B091347FD90B61AE.

Holmes:2023:BRJ

[Hol23b] Matthew Holmes. Book review:

Johan Alfredo Linthorst, Research between Science, Society and Politics: The History and Scientific Development of Green Chemistry. Utrecht: Eburon, 2023. Pp. 268. ISBN 978-94-6301-434-2. EUR36.00 (paperback). British Journal for the History of Science, 56(4):589–591, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
johan-alfredo-linthorst-
research-between-science-
society-and-politics-the-
history-and-scientific-
development-of-green-
chemistry-utrecht-eburon-
2023-pp-268-isbn-9789463014342-
3600-paperback/1FA60F7CAD4769251F0FC50153C273A5.

Holmes:2023:PWT

[Hol23c] Matthew Holmes. A plague of weasels and ticks: animal introduction, ecological disaster, and the balance of nature in Jamaica, 1870–1900. British Journal for the History of Science, 56(3):391– 407, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/plague-of-
weasels-and-ticks-animal-

                                        introduction-ecological-
                                        disaster-and-the-balance-of-
                                        nature-in-jamaica-18701900/
                                        111337F9B312A6994CBE64FCEF46E963.

Hopwood:2022:BMD

[Hop22] Nick Hopwood. ‘Not birth, marriage or death, but gastrulation’: the life of a quotation in biology. British Journal for the History of Science, 55(1):1–26, March 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        not-birth-marriage-or-death-
                                        but-gastrulation-the-life-
                                        of-a-quotation-in-biology/
                                        277AA3AE743772CB4335A98836564E48.

Horder:2023:BRM

[Hor23] Tim Horder. Book review: Marco Tamborini, The Architecture of Evolution: The Science of Form in Twentieth-Century Evolutionary Biology. Pittsburgh: University of Pittsburgh Press, 2022. Pp. 283. ISBN 978-0-8229-4735-6. $455.00 (hardcover). British Journal for the History of Science, 56 (4):591–593, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/marco-
                                        tamborini-the-architecture-
                                        of-evolution-the-science-of-
                                        form-in-twentiethcentury-
                                        evolutionary-biology-
                                        pittsburgh-university-of-

pittsburgh-press-2022-pp-
283-isbn-9780822947356-45500-
hardcover/C188762767439D9D845410586685230F.

Hall:2021:DTF

[HS21] Kersten Hall and Neeraja Sankaran. DNA translated: Friedrich Miescher’s discovery of nuclein in its original context. British Journal for the History of Science, 54(1):99–107, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/dna-translated-
friedrich-mieschers-
discovery-of-nuclein-in-its-
original-context/
60A9706BE7610FFD42F805AE636670FA.

Huang:2021:HKC

[Hua21] Chi Chi Huang. ‘Hong Kong can afford a typhoon or two’: British discussions of revolving storms. British Journal for the History of Science, 54(3):327–339, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/hong-kong-can-
afford-a-typhoon-or-two-
british-discussions-of-
revolving-storms/
75D7F1587210624FAC6E84C211D9CFA4.

Hunt:2023:BRR

[Hun23] Bruce J. Hunt. Book review:

Richard Noakes, Physics and Psychics: The Occult and the Sciences in Modern Britain. Cambridge: Cambridge University Press, 2019. Pp. 419. ISBN 978-1-316-88243-

6. £90.00 (hardcover). British Journal for the History of Science, 56(3):415–416, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
richard-noakes-physics-and-
psychics-the-occult-and-the-
sciences-in-modern-britain-
cambridge-cambridge-
university-press-2019-pp-
419-isbn-9781316882436-9000-
hardcover/
6C9F3AE766ED6473C2C4EDA493255E8B.

Ishibashi:2020:CAG

[Ish20] Yuto Ishibashi. Constructing the ‘automatic’ Greenwich time system: George Biddell Airy and the telegraphic distribution of time, c. 1852–1880. British Journal for the History of Science, 53 (1):25–46, March 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
constructing-the-automatic-
greenwich-time-system-
george-biddell-airy-and-the-
telegraphic-distribution-of-
time-c18521880/
BC3BFBBA2C010C3D73F7468EEA45AF90.

Jackson:2021:BRB

[Jac21] Roland Jackson. Book review:

Bernard Lightman and Bennet Zon (eds.), Victorian Culture and the Origin of Disciplines, New York: Routledge, 2020. Pp. 336. ISBN 978-0-367-22842-2. £120.00 (hardback). British Journal for the History of Science, 54(2):248– 250, June 2021. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/bernard-
lightman-and-bennet-zon-eds-
victorian-culture-and-the-
origin-of-disciplines-new-
york-routledge-2020-pp-336-
isbn-9780367228422-12000-
hardback/
8E25BBDC5E6B0ADFB6D6B3E2147195FE.

James:2021:BRM

[Jam21] Frank A. J. L. James. Book review: Michael Wheeler, The Athenæum: More than Just Another London Club New Haven: Yale University Press, 2020. Pp. 440. ISBN 978-0-300-24677-3. £35.00 (hardback). British Journal for the History of Science, 54(3):390–391, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/michael-wheeler-the-
athenaeum-more-than-just-
another-london-club-new-

                                        haven-yale-university-press-
                                        2020-pp-440-isbn-
                                        9780300246773-3500-hardback/
                                        B3FF9F34C6B0CC6F91A4E9C197C160C0.

Johnston:2021:MOD

[Joh21] Scott Alan Johnston. Managing the observatory: discipline, order and disorder at Greenwich, 1835–1933. British Journal for the History of Science, 54(2):155–175, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/managing-the-
                                        observatory-discipline-
                                        order-and-disorder-at-
                                        greenwich-18351933/
                                        0AAEEF7AC342DC4D5AC3AAE7D2DFC2E6.

Jones:2021:BRI

[Jon21] Andrew Jones. Book review: Ido Hartogsohn, The American Trip: Set, Setting, and the Psychedelic Experience in the Twentieth Century Cambridge, MA: MIT Press, 2020. Pp. 432. ISBN: 978-0-262-53914-

2. $35.00 (paperback). British Journal for the History of Science, 54(1):109–111, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        ido-hartogsohn-the-american-
                                        trip-set-setting-and-the-
                                        psychedelic-experience-in-
                                        the-twentieth-century-

cambridge-ma-mit-press-2020-
pp-432-isbn-9780262539142-
3500-paperback/FEE939414B35F753A8F93AE80AECA12D.

Jones:2023:SSS

[Jon23] Elin Jones. Stratifying seamanship: sailors’ knowledge and the mechanical arts in eighteenthcentury Britain. British Journal for the History of Science, 56(1):45–63, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
stratifying-seamanship-
sailors-knowledge-and-the-
mechanical-arts-in-
eighteenthcentury-britain/
C5A91723E125B35035F20F27412E1909.

Jordanova:2023:BRA

[Jor23] Ludmilla Jordanova. Book review: Andrew Cunningham, ‘I Follow Aristotle’: How William Harvey Discovered the Circulation of the Blood. London: Routledge, 2022. Pp. xii + 180. ISBN 978-1-032-16223-2. £130.00 (hardback). British Journal for the History of Science, 56:273–274, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
andrew-cunningham-i-follow-
aristotle-how-william-
harvey-discovered-the-
circulation-of-the-blood-
london-routledge-2022-pp-

                                        xii-180-isbn-9871032162232-
                                        13000-hardback/
                                        00E95A84A19D6F3F5AFEFF6DBA34D699.

Kaiser:2022:DNB

[Kai22] Katja Kaiser. Duplicate networks: the Berlin botanical institutions as a ‘clearing house’ for colonial plant material, 1891–1920. British Journal for the History of Science, 55 (3):279–296, September 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/duplicate-
                                        networks-the-berlin-
                                        botanical-institutions-as-a-
                                        clearing-house-for-colonial-
                                        plant-material-18911920/
                                        FC18720ACE9E3B4DA1748D55D8971406.

Kallstrand:2022:SNC

[Käl22] Gustav Källstrand. Science by Nobel committee: decision making and norms of scientific practice in the early physics and chemistry prizes. British Journal for the History of Science, 55(2):187–205, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        science-by-nobel-committee-
                                        decision-making-and-norms-
                                        of-scientific-practice-in-
                                        the-early-physics-and-
                                        chemistry-prizes/
                                        48B80DE55848512740AE7E05A5C77717.

Kana:2020:BRS

[Kan20] Sandip Kana. Book review:

Sabine Clarke, Science at the End of Empire: Experts and the Development of the British Caribbean, 1940–62. Manchester: Manchester University Press, 2018. Pp. 224. ISBN 978-1-5261-3138-6. £80.00 (hardback). British Journal for the History of Science, 53(4):593– 594, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/sabine-
clarke-science-at-the-end-
of-empire-experts-and-the-
development-of-the-british-
caribbean-194062-manchester-
manchester-university-press-
2018-pp-224-isbn-
9781526131386-8000-hardback/90D253E51E1B9D3B351B903774A45CB6.

Katz:2020:BRR

[Kat20] Osnat Katz. Book review: Roger D. Launius, Reaching for the Moon: A Short History of the Space Race. New Haven, CT: Yale University Press, 2019. Pp. 256. ISBN 978-0-300-23046-8. $30.00 (hardback). British Journal for the History of Science, 53(4):603– 604, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/roger-d-
launius-reaching-for-the-
moon-a-short-history-of-the-

                                        space-race-new-haven-ct-
                                        yale-university-press-2019-
                                        pp-256-isbn-9780300230468-
                                        3000-hardback/
                                        F10517977303F5486CAE4278316021B7.

Kempton:2023:CTP

[Kem23] Miles Kempton. Commercial television and primate ethology: facial expressions between Granada and London Zoo. British Journal for the History of Science, 56(1):83–102, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        commercial-television-and-
                                        primate-ethology-facial-
                                        expressions-between-granada-
                                        and-london-zoo/
                                        FEF5D5661892E52B718B001396B190C9.

Kern:2020:AEA

[Ker20] Emily M. Kern. Archaeology enters the ‘atomic age’: a short history of radiocarbon, 1946–1960. British Journal for the History of Science, 53(2):207–227, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        archaeology-enters-the-
                                        atomic-age-a-short-history-
                                        of-radiocarbon-19461960/
                                        855D6E2E66D09F7467C6051C77C2D6B0.

Khramov:2024:FDS

[Khr24] Alexander V. Khramov. A forerunner of Darwin in the service of nihilists: the translation and reception of Vestiges in Russia. British Journal for the History of Science, 57(1):65–79, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
forerunner-of-darwin-in-the-
service-of-nihilists-the-
translation-and-reception-
of-vestiges-in-russia/
891ACFF6C6BB9C45FB635D9B15C6DDCF.

Kirwan:2020:SRM

[Kir20] Adrian Kirwan. Sociability, radium and the maintenance of scientific culture and authority in twentieth-century Ireland: a case study of the Royal Dublin Society. British Journal for the History of Science, 53(1):47–66, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
sociability-radium-and-the-
maintenance-of-scientific-
culture-and-authority-in-
twentiethcentury-ireland-a-
case-study-of-the-royal-
dublin-society/
F72934AB17A85C3BD5E533EDD1D6DF55.

Khayat:2022:MAL

[KK22] Nicole Khayat and Liat Kozma.

Medicine and Arabic literary production in the Ottoman Empire during the nineteenth century.

British Journal for the History of Science, 55(4):515–524, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/medicine-and-arabic-
                                        literary-production-in-the-
                                        ottoman-empire-during-the-
                                        nineteenth-century/
                                        9EDBADC6DFE673E6B47C37672ABD4E0F.

Klerk:2020:NHP

[Kle20]  Saskia Klerk.  Natural his-

tory in the physician’s study: Jan Swammerdam (1637–1680), Steven Blankaart (1650–1705) and the ‘paperwork’ of observing insects. British Journal for the History of Science, 53(4):497–525, December 2020. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/natural-
                                        history-in-the-physicians-
                                        study-jan-swammerdam-
                                        16371680-steven-blankaart-
                                        16501705-and-the-paperwork-
                                        of-observing-insects/
                                        8D7EA3137DC5F069D32F98F38BA08896.

Kohlstedt:2023:BRH

[Koh23] Sally Gregory Kohlstedt. Book review: Hannah Wills, Sadie Harrison, Erika Jones, Rebecca Martin and Farrah Lawrence-Mackey (eds.), Women in the History of Science: A Sourcebook. London: UCL Press, 2023. Pp. xxviii + 446.

ISBN 978-1-80008-415-5. £50.00 (hardback); £30.00 (paperback); £0.00 (open-access PDF). British Journal for the History of Science, 56(4):593–595, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
hannah-wills-sadie-harrison-
erika-jones-rebecca-martin-
and-farrah-lawrencemackey-
eds-women-in-the-history-of-
science-a-sourcebook-london-
ucl-press-2023-pp-xxviii-
446-isbn-9781800084155-5000-
hardback-3000-paperback-000-openaccess-
pdf/7BBA7A0D78540A181E0FCDF044CE7805.

Komel:2024:PIS

[Kom24] Svit Komel. Petty’s instruments:

the Down Survey, territorial natural history and the birth of statistics. British Journal for the History of Science, 57(1):43–64, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/pettys-instruments-
the-down-survey-territorial-
natural-history-and-the-
birth-of-statistics/
0C35B4A6C753DE4DF81B523B4AD6B62B.

Korolija:2021:BRK

[Kor21] Maja Korolija. Book review:

Keith Ewing, Joan Mahoney and Andrew Moretta, MI5, the Cold War, and the Rule of Law. Oxford:

Oxford University Press, 2020. Pp. 528. ISBN: 978-0-19-881862-5. £84.00 (hardback). British Journal for the History of Science, 54 (4):534–536, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/keith-ewing-
                                        joan-mahoney-and-andrew-
                                        moretta-mi5-the-cold-war-
                                        and-the-rule-of-law-oxford-
                                        oxford-university-press-
                                        2020-pp-528-isbn-
                                        9780198818625-8400-hardback/
                                        6EFA68E8DABFC6F6722B1C5484DE1993.

Kotsou:2023:FIC

[Kot23] Georgiana Kotsou. Functional informality: crafting social interaction toward scientific productivity at the Gordon Research Conferences, 1950–1980. British Journal for the History of Science, 56 (4):519–534, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/functional-
                                        informality-crafting-social-
                                        interaction-toward-
                                        scientific-productivity-at-
                                        the-gordon-research-
                                        conferences-19501980/
                                        533749102A39D7A7EA41DDA0F5596E7D.

Krige:2023:BRG

[Kri23] John Krige. Book review: Gordon Barrett, China’s Cold War Science Diplomacy. Cambridge:

Cambridge University Press, 2022. Pp. 300. ISBN 978-1-1088-4457-4. £75.00 (hardback). British Journal for the History of Science, 56 (1):119–121, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/gordon-
barrett-chinas-cold-war-
science-diplomacy-cambridge-
cambridge-university-press-
2022-pp-300-isbn-
9781108844574-7500-hardback/
8B2858DB0642844D0C474798942D3DED.

Kaalund:2021:ALM

[KW21] Nanna Katrine Lüders Kaalund and John Woitkowitz. ‘Ancient lore with modern appliances’: networks, expertise, and the making of the Open Polar Sea, 1851–1853. British Journal for the History of Science, 54(3):277–299, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/ancient-lore-with-
modern-appliances-networks-
expertise-and-the-making-of-
the-open-polar-sea-18511853/
9FB26207EDAE5DBD31EE96D5EC829B0B.

Langstaff:2022:BRE

[Lan22] Alex Langstaff. Book review:

Elena Aronova, Scientific History: Experiments in History and Politics from the Bolshevik Revolution to the End of the Cold War, Chicago: The University of Chicago Press, 2021. Pp. 256. ISBN 978-0-226-76138-1. $45.00 (hardback). British Journal for the History of Science, 55(1):124– 126, March 2022. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/elena-
aronova-scientific-history-
experiments-in-history-and-
politics-from-the-bolshevik-
revolution-to-the-end-of-
the-cold-war-chicago-the-
university-of-chicago-press-
2021-pp-256-isbn-9780226761381-
4500-hardback/B1B940818A69D19F72FCD39BB374A060.

Law:2021:BRD

[Law21a] Harry Law. Book review: Dylan Mulvin, Proxies: The Cultural Work of Standing In London: MIT Press, 2021. Pp. 228. ISBN 978-0-262-04514-8. £40.00 (paperback). British Journal for the History of Science, 54(4): 532–534, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
dylan-mulvin-proxies-the-
cultural-work-of-standing-
in-london-mit-press-2021-pp-
228-isbn-9780262045148-4000-
paperback/
E01DA7CD7A6D033918C93B89506716C9.

Lawrence:2021:RMY

[Law21b] Christopher Lawrence. Robert M.

Young’s mind, brain and adaptation revisited. British Journal for the History of Science, 54(1):61–77, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
robert-m-youngs-mind-brain-
and-adaptation-revisited/
6896719470FBB886C23398325DA81035.

Law:2023:BRD

[Law23a] Harry Law. Book review: Dylan Mulvin, Proxies: The Cultural Work of Standing In London: MIT Press, 2021. Pp. 228. ISBN 978-0-262-04514-8. £40.00 (paperback). - Corrigendum. British Journal for the History of Science, 56:281, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
dylan-mulvin-proxies-the-
cultural-work-of-standing-
in-london-mit-press-2021-pp-
228-isbn-9780262045148-4000-
paperback-corrigendum/
7C4031DC2A1A212EB4EA8111953FCBD0.

Laws:2023:BRD

[Law23b] Harry Laws. Book review: Dylan Mulvin, Proxies: The Cultural Work of Standing In London: MIT Press, 2021. Pp. 228.

ISBN 978-0-2620-4514-8. £40.00 (paperback). British Journal for the History of Science, 56 (1):115–117, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
dylan-mulvin-proxies-the-
cultural-work-of-standing-
in-london-mit-press-2021-pp-
228-isbn-9780262045148-4000-
paperback/
21FD56578E1074C5F6BEA6BD0E82987D.

Lidwell-Durnin:2020:CFD

[LD20] John Lidwell-Durnin. Cultivating famine: data, experimentation and food security, 1795– 1848. British Journal for the History of Science, 53(2):159–181, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/cultivating-famine-
data-experimentation-and-
food-security-17951848/
A7FD0427A16C56306914F54969948331.

Lidwell-Durnin:2021:BRR

[LD21] John Lidwell-Durnin. Book review: Rebecca Earle, Feeding the People: The Politics of the Potato, Cambridge: Cambridge University Press, 2020. Pp. 308. ISBN 978-1-108-68845-1. £17.99 (hardback). British Journal for the History of Science, 54(2):242–244, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
rebecca-earle-feeding-the-
people-the-politics-of-the-
potato-cambridge-cambridge-
university-press-2020-pp-
308-isbn-9781108688451-1799-
hardback/
EEB700F5AAA9275294E5217CC92B0BD7.

Ledgerwood:2021:BRC

[Led21] Emmeline Ledgerwood. Book review: C. Bruce Tarter, The American Lab: An Insider’s History of the Lawrence Livermore National Laboratory, Baltimore: Johns Hopkins University Press, 2018. Pp. 453. ISBN 978-1-4214-2531-3. $79.95 (hardback). British Journal for the History of Science, 54(2):246–248, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
c-bruce-tarter-the-american-
lab-an-insiders-history-of-
the-lawrence-livermore-
national-laboratory-
baltimore-johns-hopkins-
university-press-2018-pp-
453-isbn-9781421425313-7995-
hardback/CAB848C2FE28EE1C2B1982C0E2D2ECAC.

Ledgerwood:2022:ANB

[Led22] Emmeline Ledgerwood. ‘Armed with the necessary background of knowledge’: embedding science scrutiny mechanisms in the UK Parliament. British Journal for the History of Science, 55(2):167– 185, June 2022. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/armed-with-
                                        the-necessary-background-of-
                                        knowledge-embedding-science-
                                        scrutiny-mechanisms-in-the-
                                        uk-parliament/
                                        4A0A4BB7B04D5AC31EBA8E69AEDC3B5A.

Lehnen:2023:BRM

[Leh23] Odile Manon Lehnen. Book review: Michelle DiMeo, Lady Ranelagh: The Incomparable Life of Robert Boyle’s Sister. Chicago: University of Chicago Press, 2021. Pp. 296. ISBN 978-0-226-73160-

5. $45.00 (cloth). British Journal for the History of Science, 56 (4):595–596, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/michelle-
                                        dimeo-lady-ranelagh-the-
                                        incomparable-life-of-robert-
                                        boyles-sister-chicago-
                                        university-of-chicago-press-
                                        2021-pp-296-isbn-
                                        9780226731605-4500-cloth/
                                        2495BC5A2BDB810CD566A9AFFDED2C1B.

Long:2020:CBN

[Lon20] Max Long. The ciné-biologists: natural history film and the coproduction of knowledge in interwar Britain. British Journal for the History of Science, 53(4): 527–551, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
cinebiologists-natural-
history-film-and-the-
coproduction-of-knowledge-
in-interwar-britain/
ED1CAE4195B4C3030705CD8A1D80B5AC.

Lorek:2024:BRM

[Lor24] Timothy Lorek. Book review:

Marci R. Baranski, The Globalization of Wheat, Pittsburgh: University of Pittsburgh Press, 2022. Pp. 256. ISBN 978-0-8229-4734-

9. $55.00 (hardcover). British Journal for the History of Science, 57(1):140–142, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
marci-r-baranski-the-
globalization-of-wheat-
pittsburgh-university-of-
pittsburgh-press-2022-pp-
256-isbn-9780822947349-5500-
hardcover/
66346F2CF31E50C08200E002552F1CE9.

Lowe:2021:APH

[Low21] James W. E. Lowe. Adjusting to precarity: how and why the Roslin Institute forged a leading role for itself in international networks of pig genomics research. British Journal for the History of Science, 54(4):507–530, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/adjusting-to-
                                        precarity-how-and-why-the-
                                        roslin-institute-forged-a-
                                        leading-role-for-itself-in-
                                        international-networks-of-
                                        pig-genomics-research/
                                        A4B4243748211FCEAAA55362393906EA.

Luciano:2023:ASA

[LZ23] Eugenio Luciano and Elena Zanoni. Antonio Stoppani’s ‘Anthropozoic’ in the context of the Anthropocene. British Journal for the History of Science, 56(1):103–114, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/antonio-stoppanis-
                                        anthropozoic-in-the-context-
                                        of-the-anthropocene/
                                        1E6A458F79F7028D603D68BCC8F1086B.

Luk:2024:BRJ

[LZ24] Christine Y. L. Luk and Longkai Zang. Book review: Jeffrey Womack, Radiation Evangelists: Technology, Therapy, and Uncertainty at the Turn of the Century, Pittsburgh: University of Pittsburgh Press, 2020. Pp. 288. ISBN 978-0-8229-4609-0. $35.00 (hardcover). British Journal for the History of Science, 57(1):123–125, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/jeffrey-womack-
radiation-evangelists-
technology-therapy-and-
uncertainty-at-the-turn-of-
the-century-pittsburgh-
university-of-pittsburgh-
press-2020-pp-288-isbn-
9780822946090-3500-hardcover/
35E3643D68F72CAE27C19E57C43276DE.

MacDonald:2020:BRP

[Mac20] Lee T. MacDonald. Book review:

Peter R. Broughton, Northern Star: J. S. Plaskett. Toronto and London: University of Toronto Press, 2018. Pp. xx + 539. ISBN 978-1-4426-3017-8. $90.00 (hardback). British Journal for the History of Science, 53(1):125–126, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/peter-r-broughton-
northern-star-js-plaskett-
toronto-and-london-
university-of-toronto-press-
2018-pp-xx-539-isbn-
9781442630178-9000-hardback/
D3E7D01CDD7A34B05F0267A2B7CB3AE2.

MacKinney:2022:DUH

[Mac22] Anne Greenwood MacKinney. Duplicates under the hammer: natural-history auctions in Berlin’s early nineteenth-century collection landscape. British Journal for the History of Science, 55(3):319–339, September 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
duplicates-under-the-hammer-
naturalhistory-auctions-in-
berlins-early-
nineteenthcentury-
collection-landscape/
1BED5F7051264DCD6CBA0AC7CA750C25.

Manias:2024:BRI

[Man24] Chris Manias. Book reviews: Irina Podgorny, Florentino Ameghino y hermanos, Buenos Aires: Edhasa, 2021. Pp. 346. ISBN 978-987-628-598-8. AR$9,150.00 (softcover) — Irina Podgorny, Los Argentinos vienen de los peces: Ensayo de filogenia nacional, Rosario: Beatriz Viterbo, 2021. Pp. 184. ISBN 978-9500845-399-0. AR$3,978.00 (softcover). British Journal for the History of Science, 57(1):142–144, March 2024. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
irina-podgorny-florentino-
ameghino-y-hermanos-buenos-
aires-edhasa-2021-pp-346-
isbn-9789876285988-ar915000-
softcover-irina-podgorny-
los-argentinos-vienen-de-
los-peces-ensayo-de-filogenia-
nacional-rosario-beatriz-viterbo-

2021-pp-184-isbn-97895008453990-
ar397800-softcover/99D7B6D9B2B6DBE0087F944F1B271BCE.

Marschall:2020:BRC

[Mar20] Wythe Marschall. Book review:

Chris Thorogood and Simon Hiscock, The Botany of Gin. Oxford: Bodleian Library, University of Oxford, 2020. Pp. 102. ISBN 978-1-85124-553-6. £15.00 (hardback). British Journal for the History of Science, 53(4):600–602, December 2020. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/chris-
thorogood-and-simon-hiscock-
the-botany-of-gin-oxford-
bodleian-library-university-
of-oxford-2020-pp-102-isbn-
9781851245536-1500-hardback/
190D9E8C30DBC5AE63F61BB5F3D9EDB1.

Margocsy:2022:MMG

[Mar22] Dániel Margócsy. Malinowski and malacology: global value systems and the issue of duplicates. British Journal for the History of Science, 55(3):389–409, September 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
malinowski-and-malacology-
global-value-systems-and-
the-issue-of-duplicates/
9727A997935430F98BF396C86C98718A.

Mattes:2022:MIK

[Mat22] Johannes Mattes. Mapping the invisible: knowledge, credibility and visions of earth in early modern cave maps. British Journal for the History of Science, 55(1):53–80, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        mapping-the-invisible-
                                        knowledge-credibility-and-
                                        visions-of-earth-in-early-
                                        modern-cave-maps/
                                        9A626594DED16EB9785EDA54E150EF2B.

Mauuarin:2022:VDS

[Mau22] Anaïs Mauuarin. Visual duplication: specimens, works of art and photographs at the Musée d’ethnographie du Trocadéro (1928–1935). British Journal for the History of Science, 55 (3):365–388, September 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/visual-
                                        duplication-specimens-works-
                                        of-art-and-photographs-at-
                                        the-musee-dethnographie-du-
                                        trocadero-19281935/
                                        0618FC4683D60D7B595E7424D9B36C34.

Murphy:2022:REG

[MB22] Jane H. Murphy and Sahar Bazzaz. Re-examining globalization and the history of science: Ottoman and Middle Eastern experiences. British Journal for the History of Science, 55(4): 411–422, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
reexamining-globalization-
and-the-history-of-science-
ottoman-and-middle-eastern-
experiences/
7EE29C456A52DD9DAF5417C94AD82615.

McAleer:2022:TCW

[McA22] John McAleer. ‘The troubles of collecting’: William Henry Harvey and the practicalities of naturalhistory collecting in Britain’s nineteenth-century world. British Journal for the History of Science, 55(1):81–100, March 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
troubles-of-collecting-
william-henry-harvey-and-
the-practicalities-of-
naturalhistory-collecting-
in-britains-
nineteenthcentury-world/
80BD873A0B1A3B2B49C595F12322C7E1.

McCahey:2023:BRN

[McC23] Daniella McCahey. Book review: Nanna Katrine Lüders Kaalund, Explorations in the Icy North: How Travel Narratives Shaped Arctic Science in the Nineteenth Century. Pittsburgh: University of Pittsburgh Press, 2021. Pp. 230. ISBN 978-0-8229-4659-5. $40.00 (hardback). British Journal for the History of Science, 56 (1):123–124, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/nanna-
                                        katrine-luders-kaalund-
                                        explorations-in-the-icy-
                                        north-how-travel-narratives-
                                        shaped-arctic-science-in-
                                        the-nineteenth-century-
                                        pittsburgh-university-of-
                                        pittsburgh-press-2021-pp-230-
                                        isbn-9780822946595-4000-hardback/
                                        836F6A8B87B59DC1C0A7EF6A5AF90141.

McGovern:2021:GGD

[McG21] Michael F. McGovern. Genes go digital: Mendelian Inheritance in Man and the genealogy of electronic publishing in biomedicine. British Journal for the History of Science, 54(2):213–231, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        genes-go-digital-mendelian-
                                        inheritance-in-man-and-the-
                                        genealogy-of-electronic-
                                        publishing-in-biomedicine/
                                        CFB5307BA6645D9FA9A9634B03DB0F4B.

McMullan:2020:BRS

[McM20] Joshua McMullan. Book review: Sergei Belyakov, Liquidator: The Chernobyl Story. Singapore: World Scientific Publishing, 2019. Pp. 188. ISBN 978-981-322-868-9. £25.00 (paperback). British Journal for the History of Science, 53(4):599–600, December 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/sergei-belyakov-
liquidator-the-chernobyl-
story-singapore-world-
scientific-publishing-2019-
pp-188-isbn-9789813228689-
2500-paperback/
7378AD3866DDADC45DF7C378170BE252.

Medori:2023:VDC

[Med23] Beatriz Medori. The visual diplomacy of cancer treatments: the mediatic legacy of the Curies in the early transnational fight against cancer. British Journal for the History of Science, 56:167–183, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
visual-diplomacy-of-cancer-
treatments-the-mediatic-
legacy-of-the-curies-in-the-
early-transnational-fight-
against-cancer/
19E901018D2329B8BFBF6A34B150171C.

Meyer:2020:BRH

[Mey20] Tirza Meyer. Book review: Helen M. Rozwadowski, Vast Expanses: A History of the Oceans. London: Reaktion Books, 2018. Pp. 268. ISBN 978-1-78023-997-

2. £16.00 (hardback). British Journal for the History of Science, 53(1):123–125, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        helen-m-rozwadowski-vast-
                                        expanses-a-history-of-the-
                                        oceans-london-reaktion-
                                        books-2018-pp-268-isbn-
                                        9781780239972-1600-hardback/
                                        C09B1BC3A65C3AE2607486BB62CE3C93.

Meyer:2022:ILS

[Mey22] Victoria N. Meyer. Innovations from the Levant: smallpox inoculation and perceptions of scientific medicine. British Journal for the History of Science, 55(4): 423–444, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/innovations-
                                        from-the-levant-smallpox-
                                        inoculation-and-perceptions-
                                        of-scientific-medicine/
                                        E3C38F05248CF4393EA2EFD26D743536.

Middleton:2021:BRE

[Mid21] Lindsay Middleton. Book review: Emma Griffin, Bread Winner: An Intimate History of the Victorian Economy. New Haven, CT: Yale University Press, 2020. Pp. 320. ISBN 978-0-300-23006-2. £20.00 (hardback). British Journal for the History of Science, 54(1):117–118, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
emma-griffin-bread-winner-
an-intimate-history-of-the-
victorian-economy-new-haven-
ct-yale-university-press-
2020-pp-320-isbn-
9780300230062-2000-hardback/
68EBFE468C62FF294000445E3679235B.

Miller:2021:BRA

[Mil21] Ian Miller. Book review: Andrew B. Liu, Tea War: A History of Capitalism in China and India. New Haven, CT, and London: Yale University Press, 2020. Pp. 360. ISBN 978-0-300-24373-

4. £35.00 (hardcover). British Journal for the History of Science, 54(3):391–392, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
andrew-b-liu-tea-war-a-
history-of-capitalism-in-
china-and-india-new-haven-
ct-and-london-yale-
university-press-2020-pp-
360-isbn-9780300243734-3500-
hardcover/
286A1F6141DC7781B74BD48ED874D5D2.

Morris:2021:BRT

[Mor21a] Andrew M. A. Morris. Book review: Toby Musgrave, The Multifarious Mr Banks: From Botany Bay to Kew, the Natural Historian Who Shaped the World, New Haven, CT: Yale University Press, 2020. Pp. 386. ISBN 978-0-300-22383-5. $35.00 (hardcover). British Journal for the History of Science, 54(2):245–246, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/toby-musgrave-the-
                                        multifarious-mr-banks-from-
                                        botany-bay-to-kew-the-
                                        natural-historian-who-
                                        shaped-the-world-new-haven-
                                        ct-yale-university-press-
                                        2020-pp-386-isbn-
                                        9780300223835-3500-hardcover/
                                        2CE89541304076C7522A1B94E23DCB7B.

Morus:2021:FWH

[Mor21b] Iwan Rhys Morus. Out on the fringe: Wales and the history of science. British Journal for the History of Science, 54(1):87–97, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        out-on-the-fringe-wales-and-
                                        the-history-of-science/
                                        69B7B8BEB81ACCA9796117210D8EB5C0.

Mougey:2023:NNI

[Mou23] Thomas Mougey. Negotiating the norms of an international science: standardization work at the International Geological Congress, 1878–1891. British Journal for the History of Science, 56(4): 435–451, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
negotiating-the-norms-of-an-
international-science-
standardization-work-at-the-
international-geological-
congress-18781891/
73ADE4A8ADCA0671C8106A816D19E100.

Moxham:2021:BRM

[Mox21] Noah Moxham. Book review:

Maria M. Portuondo, The Spanish Disquiet: The Biblical Natural Philosophy of Benito Arias Montano. Chicago and London: The University of Chicago Press, 2019. Pp. xvi + 446. ISBN 978-0-226-59226-8. $65.00 (hardback). British Journal for the History of Science, 54(1):126–128, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/maria-m-portuondo-
the-spanish-disquiet-the-
biblical-natural-philosophy-
of-benito-arias-montano-
chicago-and-london-the-
university-of-chicago-press-
2019-pp-xvi-446-isbn-
9780226592268-6500-hardback/E80CFCA92FA68434AA51992BFA6C36EB.

Michel:2021:MAD

[MS21] Nicolas Michel and Ivahn Smadja. Mathematics in the archives: deconstructive historiography and the shaping of modern geometry (1837–1852). British Journal for the History of Science, 54(4): 423–441, December 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        mathematics-in-the-archives-
                                        deconstructive-
                                        historiography-and-the-
                                        shaping-of-modern-geometry-
                                        18371852/
                                        8C08FCC4DEE5B9C2C846D072808C727F.

Muller:2024:BRO

[Mül24] Janna K. Müller. Book review: Omar W. Nasim, The Astronomer’s Chair: A Visual and Cultural History, Cambridge, MA: MIT Press, 2021. Pp. 312. ISBN 978-0-262-04553-7. $60.00 (paperback). British Journal for the History of Science, 57(1):131–132, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/omar-w-nasim-the-
                                        astronomers-chair-a-visual-
                                        and-cultural-history-
                                        cambridge-ma-mit-press-2021-
                                        pp-312-isbn-9780262045537-
                                        6000-paperback/
                                        95A8C2A09B8CA466A70DE2DB953026E2.

Niala:2024:BRA

[Nia24] J. C. Niala. Book review: Anna Marie Roos and Vera Keller (eds), Collective Wisdom: Collecting in the Early Modern Academy, Turnhout: Brepols, 2022. Pp. 325. ISBN 978-2-503-58806-3. EUR85.00 (hardback). British Journal for the History of Science, 57(1):136–138, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/anna-marie-roos-and-
vera-keller-eds-collective-
wisdom-collecting-in-the-
early-modern-academy-
turnhout-brepols-2022-pp-
325-isbn-9782503588063-8500-
hardback/4D8443581C2EB4C6BE5520A1536718F7.

Nichols:2022:CDO

[Nic22] Catherine A. Nichols. Curating duplicates: operationalizing similarity in the Smithsonian Institution with Haida rattles, 1880– 1926. British Journal for the History of Science, 55(3):341–363, September 2022. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/curating-
duplicates-operationalizing-
similiarity-in-the-
smithsonian-institution-
with-haida-rattles-18801926/
845161923F4D901FB9388C4E723A2435.

Nixon:2024:WBO

[Nix24] Sean Nixon. Watching birds: observation, photography and the ‘ethological eye’. British Journal for the History of Science, 57(1):1–19, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        watching-birds-observation-
                                        photography-and-the-
                                        ethological-eye/
                                        BCFB24B7609DFB5A113C34F5D46E87ED.

Olshin:2023:BRA

[Ols23] Benjamin B. Olshin. Book review: Alida C. Metcalf, Mapping an Atlantic World, circa 1500. Baltimore: Johns Hopkins University Press, 2020. Pp. 224. ISBN 978-1-4214-3852-8. $57.00 (hardback). British Journal for the History of Science, 56(4):597–598, December 2023. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/alida-c-
                                        metcalf-mapping-an-atlantic-
                                        world-circa-1500-baltimore-
                                        johns-hopkins-university-
                                        press-2020-pp-224-isbn-
                                        9781421438528-5700-hardback/
                                        4A5C1785AF591FEB4F061F64FB3704AB.

Palmieri:2021:BRA

[Pal21] Kristine Palmieri. Book reviews: Alex Csiszar, The Scientific Journal: Authorship and the Politics of Knowledge in the Nineteenth Century Cambridge, MA: Harvard University Press, 2018. Pp. 368. ISBN 978-0-226-55323-

8. $45.00 (hardback). ISBN 978-0-226-75250-1. $35.00 (paperback). British Journal for the History of Science, 54(2):241–242, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
alex-csiszar-the-scientific-
journal-authorship-and-the-
politics-of-knowledge-in-
the-nineteenth-century-
cambridge-ma-harvard-
university-press-2018-pp-
368-isbn-9780226553238-4500-
hardback-isbn-9780226752501-3500-
paperback/EED2D4E3ED6727E73F7114024795DC1D.

Pfeffer:2021:SAC

[Pfe21] Michelle Pfeffer. The Society of Astrologers (c. 1647–1684): sermons, feasts and the resuscitation of astrology in seventeenthcentury London. British Journal for the History of Science, 54(2):133–153, June 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
society-of-astrologers-
c16471684-sermons-feasts-
and-the-resuscitation-of-
astrology-in-

                                        seventeenthcentury-london/
                                        B292AB39DACD6474CBAB1000BAE4E655.

Piel:2020:SBS

[Pie20] Helen Piel. Scientific broadcasting as a social responsibility? John Maynard Smith on radio and television in the 1960s and 1970s. British Journal for the History of Science, 53(1):89–108, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        scientific-broadcasting-as-
                                        a-social-responsibility-
                                        john-maynard-smith-on-radio-
                                        and-television-in-the-1960s-
                                        and-1970s/
                                        6F06499A7EDA8C0282E56DB79BA95F9A.

Ranford:2023:BRA

[Ran23] Paul Ranford. Book review: Aileen Fyfe, Noah Moxham, Julie McDougal-Waters and Camilla Mørk Røstvik, A History of Scientific Journals: Publishing at the Royal Society, 1665–2015. London: UCL Press, 2022. Pp. 643. ISBN 978-1-8000-8234-2. £60.00 (hardcover), £0.00 (open-access pdf). Doi:10.14324/111.9781800082328. British Journal for the History of Science, 56(3):416–418, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/aileen-fyfe-noah-

moxham-julie-mcdougalwaters-
and-camilla-mork-rostvik-a-
history-of-scientific-
journals-publishing-at-the-royal-
society-16652015-london-ucl-press-
2022-pp-643-isbn-9781800082342-
6000-hardcover-000-openaccess-
pdf-doi10143241119781800082328/
0C60483D34C81972D3188340AB6D81AD.

Robinson:2020:ICR

[RBB^{+}20] Sam Robinson, Megan Baumhammer, Lea Beiermann, Daniel Belteki, Amy C. Chambers, Kelcey Gibbons, Edward Guimont, Kathryn Heffner, Emma-Louise Hill, Jemma Houghton, Daniella Mccahey, Sarah Qidwai, Charlotte Sleigh, Nicola Sugden, and James Sumner. Innovation in a crisis: rethinking conferences and scholarship in a pandemic and climate emergency. British Journal for the History of Science, 53(4): 575–590, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
innovation-in-a-crisis-
rethinking-conferences-and-
scholarship-in-a-pandemic-
and-climate-emergency/
601948CB0B684F833AC3E96643E46F2C.

Reeves:2021:BRT

[Ree21] Carole Reeves. Book review:

Tom Sharpe, The Fossil Woman: A Life of Mary Anning Wimborne, Minster: The Dovecote Press, 2020. Pp. 240. ISBN 978-0-9955462-9-5. £20.00 (hardback).

British Journal for the History of Science, 54(2):250–251, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        tom-sharpe-the-fossil-woman-
                                        a-life-of-mary-anning-
                                        wimborne-minster-the-
                                        dovecote-press-2020-pp-240-
                                        isbn-9780995546295-2000-
                                        hardback/
                                        3095C57F79274F868CC99B7AD6BFC2BF.

Reinisch:2023:TCT

[Rei23] Jessica Reinisch. Technical conferences as a technique of internationalism. British Journal for the History of Science, 56(4):485– 502, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/technical-
                                        conferences-as-a-technique-
                                        of-internationalism/
                                        72E819978160AF024A4D37CFA7C3A83F.

Rispoli:2020:BRN

[Ris20] Giulia Rispoli. Book review: Nicholas B. Breyfogle (ed.), Eurasian Environments: Nature and Ecology in Imperial Russian and Soviet History. Pittsburgh: University of Pittsburgh Press, 2018. Pp. 424. ISBN: 978-0-8229-6563-

3. $34.95 (paperback). British Journal for the History of Science, 53(1):134–136, March 2020.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
nicholas-b-breyfogle-ed-
eurasian-environments-
nature-and-ecology-in-
imperial-russian-and-soviet-
history-pittsburgh-
university-of-pittsburgh-
press-2018-pp-424-isbn-9780822965633-
3495-paperback/3560B54516BF0D92C6EC2C7844BF9390.

Rossini:2020:BRR

[Ros20] Paolo Rossini. Book review:

Richard J. Oosterhoff, Making Mathematical Culture: University and Print in the Circle of Lefèvre d’Etaples, Oxford: Oxford University Press, 2018. Pp. 304. ISBN 978-0-19-882352-0. £65.00 (hardback). British Journal for the History of Science, 53(2):282– 283, June 2020. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
richard-j-oosterhoff-making-
mathematical-culture-
university-and-print-in-the-
circle-of-lefevre-detaples-
oxford-oxford-university-
press-2018-pp-304-isbn-
9780198823520-6500-hardback/03675126FF50009A9FA3E88FDFEDF381.

Rose:2022:BRT

[Ros22] Sebastian James Rose. Book review: Thomas Simpson, The Frontier in British India: Space, Science and Power in the Nineteenth Century. Cambridge: Cambridge University Press, 2021. Pp. xvi + 298. ISBN 978-1-108-87915-6. £75.00 (hardback). British Journal for the History of Science, 55 (2):250–252, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/thomas-
                                        simpson-the-frontier-in-
                                        british-india-space-science-
                                        and-power-in-the-nineteenth-
                                        century-cambridge-cambridge-
                                        university-press-2021-pp-
                                        xvi-298-isbn-9781108879156-
                                        7500-hardback/0B9213454159BD3937E12A3324C28217.

Rossetter:2023:AVN

[Ros23] Thomas Rossetter. Anti-voluntarism, natural providence and miracles in Thomas Burnet’s theory of the Earth. British Journal for the History of Science, 56(1):1–20, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/antivoluntarism-
                                        natural-providence-and-
                                        miracles-in-thomas-burnets-
                                        theory-of-the-earth/
                                        029A224C155D5BCBD774177896007B40.

Sachdev:2021:BRD

[Sac21] Vorathep Sachdev. Book review: David Herzberg, White Market Drugs: Big Pharma and the Hidden History of Addiction in America, Chicago: The University of Chicago Press, 2020. ISBN 978-0-226-73188-9. $27.50 (hardback). British Journal for the History of Science, 54(2):252–253, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
david-herzberg-white-market-
drugs-big-pharma-and-the-
hidden-history-of-addiction-
in-america-chicago-the-
university-of-chicago-press-
2020-isbn-9780226731889-
2750-hardback/
172E9835A008956F1E5C2FDE05377EA8.

Schleper:2022:CCT

[Sch22a] Simone Schleper. Caribou crossings: the Trans-Alaska Pipeline System, conservation, and stakeholdership in the Anthropocene. British Journal for the History of Science, 55(2):127–143, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
caribou-crossings-the-
transalaska-pipeline-system-
conservation-and-
stakeholdership-in-the-
anthropocene/
B1CB3827444C3D4CA70654544D59F217.

Schmechel:2022:DFD

[Sch22b] Carmen Schmechel. Descartes on fermentation in digestion: iatromechanism, analogy and teleology. British Journal for the History of Science, 55(1):101–116, March 2022. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        descartes-on-fermentation-
                                        in-digestion-iatromechanism-
                                        analogy-and-teleology/
                                        7E372D66D0B035A876A7FB92C0A0A525.

Secord:2021:RHDa

[Sec21a] James A. Secord. Revolutions in the head: Darwin, Malthus and Robert M. Young. British Journal for the History of Science, 54 (1):41–59, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

                                        https://www.cambridge.org/
                                        core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/revolutions-
                                        in-the-head-darwin-malthus-
                                        and-robert-m-young/
                                        9DBEACA1949208DD36BF77C8653DF21E.

See corrigendum [Sec21b].

Secord:2021:RHDb

[Sec21b] James A. Secord. Revolutions in the head: Darwin, Malthus and Robert M. Young — corrigendum. British Journal for the History of Science, 54(2):255, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/

article/revolutions-in-the-
head-darwin-malthus-and-
robert-m-young-corrigendum/
073D3E929F2BB8B5C21C8F62A8EFCB63.

See [Sec21a].

Secord:2022:BRD

[Sec22] James A. Secord. Book review:

Diarmid A. Finnegan, The Voice of Science: British Scientists on the Lecture Circuit in Gilded Age America, Pittsburgh: University of Pittsburgh Press, 2021. Pp. xiii + 286. ISBN 978-0-8229-4681-

6. $60.00 (hardback). British Journal for the History of Science, 55(1):117–119, March 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
diarmid-a-finnegan-the-
voice-of-science-british-
scientists-on-the-lecture-
circuit-in-gilded-age-
america-pittsburgh-
university-of-pittsburgh-
press-2021-pp-xiii-286-isbn-
9780822946816-6000-hardback/8A582F1D72E0A337B17DCB3818CDAB28.

Shalvey:2022:BRA

[Sha22] Aisling Shalvey. Book review:

Amir Teicher, Social Mendelism: Genetics and the Politics of Race in Germany, 1900–1948. Cambridge: Cambridge University Press, 2021. Pp. xiv + 268. ISBN 978-1-108-49949-1. £26.99 (hardback). British Journal for the History of Science, 55(2):252–253, June 2022. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/amir-teicher-social-
mendelism-genetics-and-the-
politics-of-race-in-germany-
19001948-cambridge-
cambridge-university-press-
2021-pp-xiv-268-isbn-
9781108499491-2699-hardback/
86E83B483AFB11EFB7174BBCC0B2BD06.

Short:2022:BRA

[Sho22] Devin Short. Book review: Alex Wellerstein, Restricted Data : The History of Nuclear Secrecy in the United States, Chicago: The University of Chicago Press, 2021. Pp. 528. ISBN 978-0-2260-2038-9. $35.00 (hardback). British Journal for the History of Science, 55 (4):525–527, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
alex-wellerstein-restricted-
data-the-history-of-nuclear-
secrecy-in-the-united-
states-chicago-the-
university-of-chicago-press-
2021-pp-528-isbn-
9780226020389-3500-hardback/
F090D1B68D646FCABA0F6BBB42A8FA06.

Sialaros:2021:BRA

[Sia21] Michalis Sialaros. Book review: Alexander Jones and Liba Taub (eds.), The Cambridge History of Science, vol. 1, Ancient Science Cambridge: Cambridge University Press, 2018. Pp. xix + 642. ISBN 978-0-511-98014-5. £120.00 (hardback). British Journal for the History of Science, 54(1):124–125, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/alexander-jones-and-
liba-taub-eds-the-cambridge-
history-of-science-vol-1-
ancient-science-cambridge-
cambridge-university-press-
2018-pp-xix-642-isbn-
9780511980145-12000-
hardback/FE05FCF78E3727110F679FC440C82CA5.

Sime:2022:BRD

[Sim22] Zane Sime. Book review: Dipesh Chakrabarty, The Climate of History in a Planetary Age. Chicago: The University of Chicago Press, 2021. Pp. 296. ISBN 978-0-226-73286-2. $25.00 (paperback). British Journal for the History of Science, 55 (2):247–248, June 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/dipesh-
chakrabarty-the-climate-of-
history-in-a-planetary-age-
chicago-the-university-of-
chicago-press-2021-pp-296-
isbn-9780226732862-2500-
paperback/18775C1057A57BA5A145286D5A3F0CFA.

Sleigh:2020:BRJ

[Sle20] Charlotte Sleigh. Book review: Joshua Nall, News from Mars: Mass Media and the Forging of a New Astronomy, 1860–1910. Pittsburgh: University of Pittsburgh Press, 2019. Pp. 287. ISBN 978-0-8229-4552-9. $50.00 (hardcover). British Journal for the History of Science, 53(2):281–282, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        joshua-nall-news-from-mars-
                                        mass-media-and-the-forging-
                                        of-a-new-astronomy-18601910-
                                        pittsburgh-university-of-
                                        pittsburgh-press-2019-pp-
                                        287-isbn-9780822945529-5000-
                                        hardcover/F1CAFB3BB21EAF146C16937F14590BB2.

Smith:2020:BRK

[Smi20] Roger Smith. Book review: Kerri Andrews, Wanderers: A History of Women Walking. London: Reaktion Books, 2020. Pp. 303. ISBN 978-1-78914-342-3. £14.99 (hardback). British Journal for the History of Science, 53(4): 602–603, December 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        kerri-andrews-wanderers-a-
                                        history-of-women-walking-
                                        london-reaktion-books-2020-

pp-303-isbn-9781789143423-
1499-hardback/
5485980149031EE578A093D9DA06771B.

Smith:2021:MPC

[Smi21a] Crosbie Smith. Maurice Pierre Crosland (1931–2020): an appreciation. British Journal for the History of Science, 54(1):79– 85, March 2021. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
maurice-pierre-crosland-
19312020-an-appreciation/
58825CA57DD603492B21B855A4410F15.

Smith:2021:BRE

[Smi21b] David F. Smith. Book review:

Elizabeth A. Williams, Appetite and Its Discontents: Science, Medicine and the Urge to Eat, 1750–1950. Chicago: The University of Chicago Press, 2020. Pp. 416. ISBN 978-0-226-69304-0. $35.00 (paperback). British Journal for the History of Science, 54 (1):128–129, March 2021. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/elizabeth-a-
williams-appetite-and-its-
discontents-science-
medicine-and-the-urge-to-
eat-17501950-chicago-the-
university-of-chicago-press-
2020-pp-416-isbn-

9780226693040-3500-paperback/
751ED7088372EA7728CF388E24CAAAA2.

Smith:2023:BRR

[Smi23] Roger Smith. Book review: Robert J. Sternberg and Wade E. Pickren (eds.), The Cambridge Handbook of the Intellectual History of Psychology, Cambridge: Cambridge University Press, 2019. Pp. 533. ISBN 978-1-108-41869-

0. $210.00 (hardback). British Journal for the History of Science, 56(3):418–419, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
robert-j-sternberg-and-wade-
e-pickren-eds-the-cambridge-
handbook-of-the-
intellectual-history-of-
psychology-cambridge-
cambridge-university-press-
2019-pp-533-isbn-9781108418690-
21000-hardback/63BC8389BE66B42918D52FC046D86545.

Smith:2024:BRM

[Smi24] Roger Smith. Book review: Margarete Vöhringer, Avant-Garde and Psychotechnics: Science, Art and Technology in the Early Soviet Union, London: Routledge, 2023. Pp. 254. ISBN 978-1-032-53264-6. £104.00 (hardback). British Journal for the History of Science, 57 (1):133–134, March 2024. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-

science/article/margarete-
vohringer-avantgarde-and-
psychotechnics-science-art-
and-technology-in-the-early-
soviet-union-london-
routledge-2023-pp-254-isbn-
9781032532646-10400-
hardback/8022A3E305194225A8747BA104FA1E95.

Somsen:2023:GWS

[Som23] Geert Somsen. ‘The goddess that we serve’: projecting international community at the first serial chemistry conferences, 1893– 1914. British Journal for the History of Science, 56(4):453–467, December 2023. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/goddess-
that-we-serve-projecting-
international-community-at-
the-first-serial-chemistry-
conferences-18931914/
53349DC11128CC0F6CA961B59D24641D.

Spinardi:2020:GAN

[Spi20] Graham Spinardi. The green airliner that never was: aerodynamic theory, fuel-efficiency and the role of the British state in aviation technology in the midtwentieth century. British Journal for the History of Science, 53(2):229–254, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/

                                        green-airliner-that-never-
                                        was-aerodynamic-theory-
                                        fuelefficiency-and-the-role-
                                        of-the-british-state-in-
                                        aviation-technology-in-the-
                                        midtwentieth-century/
                                        62F9ABAB856994206E46E389CA0FF08A.

Spielmann:2023:BRP

[Spi23] Sophia Spielmann. Book review: Paula De Vos, Compound Remedies: Galenic Pharmacy from the Ancient Mediterranean to New Spain. Pittsburgh: University of Pittsburgh Press, 2021. Pp. 404. ISBN 978-0-8229-4649-6 $50.00 (hardback). British Journal for the History of Science, 56:275–276, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/paula-de-vos-
                                        compound-remedies-galenic-
                                        pharmacy-from-the-ancient-
                                        mediterranean-to-new-spain-
                                        pittsburgh-university-of-
                                        pittsburgh-press-2021-pp-
                                        404-isbn-9780822946496-5000-
                                        hardback/B7B096A7041F9ECA92403AC92F4DBE59.

Sera-Shriar:2023:MDF

[SS23] Efram Sera-Shriar. From museumization to decolonization: fostering critical dialogues in the history of science with a Haida eagle mask. British Journal for the History of Science, 56(3):309–328, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/

journals/british-journal-
for-the-history-of-science/
article/from-museumization-
to-decolonization-fostering-
critical-dialogues-in-the-
history-of-science-with-a-
haida-eagle-mask/
E6B3D4C885685F0A1843F7A099CCE380.

Stark:2022:RDD

[Sta22] James F. Stark. ‘A remedy for this dread disease’: Achille Sclavo, anthrax and serum therapy in early twentieth-century Britain. British Journal for the History of Science, 55(2):207–226, June 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/remedy-for-
this-dread-disease-achille-
sclavo-anthrax-and-serum-
therapy-in-early-
twentiethcentury-britain/
DBFFB0865A244BB7B34C290BB08FD690.

Stenhouse:2022:BRA

[Ste22] Brigitte Stenhouse. Book reviews:

Allan Esterson and David C. Cassidy, Einstein’s Wife: The Real Story of Mileva Einstein-Marić Cambridge, MA: MIT Press, 2019. Pp. 336. ISBN: 978-0-2625-3897-

8. $19.95 (paperback). ISBN: 978-0-2620-3961-1. $29.95 (hardback). British Journal for the History of Science, 55(4):531–533, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-

                                        for-the-history-of-science/
                                        article/allan-esterson-and-
                                        david-c-cassidy-einsteins-
                                        wife-the-real-story-of-
                                        mileva-einsteinmaric-
                                        cambridge-ma-mit-press-2019-
                                        pp-336-isbn-9780262538978-
                                        1995-paperback-isbn-
                                        9780262039611-2995-hardback/EC290FBBFB3E58DC60E32E6AE4B6CD49.

Stepney:2024:VCO

[Ste24] Eva Surawy Stepney. ‘Visible’ compulsions: OCD and the politics of science in British clinical psychology, 1948–1975. British Journal for the History of Science, 57(1):81–97, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        visible-compulsions-ocd-and-
                                        the-politics-of-science-in-
                                        british-clinical-psychology-
                                        19481975/
                                        D431B7D6003860F9E6ABE50476BA46A4.

Storni:2021:DPD

[Sto21] Marco Storni. Denis Papin’s digester and its eighteenth-century European circulation. British Journal for the History of Science, 54(4):443–463, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        denis-papins-digester-and-
                                        its-eighteenthcentury-

european-circulation/
81444C8189DA86253389A007FED65612.

Stolz:2022:IPA

[Sto22] Daniel A. Stolz. ‘Impossible to provide an accurate estimate’: the interested calculation of the Ottoman public debt, 1875–1881. British Journal for the History of Science, 55(4):477–493, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/impossible-to-
provide-an-accurate-
estimate-the-interested-
calculation-of-the-ottoman-
public-debt-18751881/
9FCA5C2D392B2CB590DFF48DECA8847F.

Stoeger:2023:BRF

[Sto23] Alexander Stoeger. Book review: Felix Lüttge, Auf den Spuren des Wals: Geographien des Lebens im 19. Jahrhundert Göttingen: Wallstein Verlag, 2020. Pp. 279. ISBN 978-3-8353-3680-

3. EUR28.00 (hardback). British Journal for the History of Science, 56(1):129–131, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
felix-luttge-auf-den-spuren-
des-wals-geographien-des-
lebens-im-19-jahrhundert-
gottingen-wallstein-verlag-
2020-pp-279-isbn-

                                        9783835336803-2800-hardback/
                                        F449FCAD458A941AA4A92DBB738259AD.

Turchetti:2023:IPI

[TA23] Simone Turchetti and Matthew Adamson. Introduction: Power to the image! science, technology and visual diplomacy. British Journal for the History of Science, 56:135–146, June 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        introduction-power-to-the-
                                        image-science-technology-
                                        and-visual-diplomacy/
                                        FDE4F52D3091011E66B1068CF31FEB43.

Talairach:2021:BRC

[Tal21] Laurence Talairach. Book reviews: Charles Smith, James T. Costa and David Collard (eds.), An Alfred Russel Wallace Companion, Chicago and London: The University of Chicago Press, 2019. Pp. 416. ISBN 978-0-226-62210-

1. $60.00 (hardback). Michael A. Flannery, Nature’s Prophet: Alfred Russel Wallace and His Evolution from Natural Selection to Natural Theology, Tuscaloosa: University of Alabama Press, 2018. Pp. 280. ISBN 978-0-8173-1985-4. $44.95 (hardback). British Journal for the History of Science, 54(2):233–236, June 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-

history-of-science/article/
charles-smith-james-t-costa-
and-david-collard-eds-an-
alfred-russel-wallace-
companion-chicago-and-
london-the-university-of-
chicago-press-2019-pp-416-
isbn-9780226622101-6000-hardback-
michael-a-flannery-natures-prophet-
alfred-russel-wallace-and-his-
evolution-from-natural-selection-
to-natural-theology-tuscaloosa-
university-of-alabama-press-2018-
pp-280-isbn-9780817319854-4495-
hardback/50DEC7EDA05762B4CAB0638B3053CCAC.

Tarrant:2022:BRM

[Tar22] Neil Tarrant. Book review: Mark A. Waddell, Magic, Science and Religion in Early Modern Europe, Cambridge: Cambridge University Press, 2021. Pp. 220. ISBN 978-1-108-34823-2. £69.99/£19.99 (hardback/paperback). British Journal for the History of Science, 55(1):123–124, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL https://www.cambridge.
org/core/journals/british-
journal-for-the-history-of-
science/article/mark-a-
waddell-magic-science-and-
religion-in-early-modern-
europe-cambridge-cambridge-
university-press-2021-pp-
220-isbn-9781108348232-
69991999-hardbackpaperback/D67F2CAB7EC534A85DC4CCBBA60755AE.

Tarrant:2024:BRM

[Tar24] Neil Tarrant. Book review:

Mackenzie Cooley, The Perfection of Nature: Animals, Breeding, and Race in the Renaissance, Chicago: University of Chicago Press, 2022. Pp. 334. ISBN 978-0-226-82228-0. $112.50 (cloth). British Journal for the History of Science, 57(1):138–140, March 2024. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/mackenzie-cooley-
                                        the-perfection-of-nature-
                                        animals-breeding-and-race-
                                        in-the-renaissance-chicago-
                                        university-of-chicago-press-
                                        2022-pp-334-isbn-
                                        9780226822280-11250-cloth/
                                        C18C2FD82F97E419C865CFC7F313D1D3.

Tekin:2022:IPG

[Tek22] Kenan Tekin. Islamic philosophy and the globalization of science: Ahmed Cevdet’s translation of the sixth chapter of Ibn Khaldun’s Muqaddimah. British Journal for the History of Science, 55 (4):459–475, December 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        islamic-philosophy-and-the-
                                        globalization-of-science-
                                        ahmed-cevdets-translation-
                                        of-the-sixth-chapter-of-ibn-
                                        khalduns-muqaddimah/
                                        C060C8FFB6B8A6E51B0DA2E6835316C9.

teHennepe:2023:VLF

[tH23]  Mieneke  te  Hennepe.

Van Leeuwenhoek — the film: remaking memory in Dutch science cinema 1925–c. 1960. British Journal for the History of Science, 56(3):329–349, September 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
van-leeuwenhoek-the-film-
remaking-memory-in-dutch-
science-cinema-1925c1960/
A9BDEFD1D3BA838EAFF76231C74EC530.

Timberlake:2020:BRM

[Tim20] Todd K. Timberlake. Book review: Martin Beech, Going Underground: The Science and History of Falling through the Earth. New Jersey, London, Singapore, Beijing, Shanghai, Hong Kong, Taipei, Chennai and Tokyo: World Scientific, 2019. Pp. xi + 276. ISBN 978-981-327-903-2. £35.00/ $38.00 (paperback). British Journal for the History of Science, 53(2):287–288, June 2020. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
martin-beech-going-
underground-the-science-and-
history-of-falling-through-
the-earth-new-jersey-london-
singapore-beijing-shanghai-
hong-kong-taipei-chennai-
and-tokyo-world-scientific-
2019-pp-xi-276-isbn-9789813279032-
35003800-paperback/FFB6C23B4A6BBD7D65A62AD52C606FE7.

Turney:2021:BRP

[Tur21] Jon Turney. Book reviews: Patrick Manning, A History of Humanity: The Evolution of the Human System. Cambridge: Cambridge University Press, 2020. Pp. xiii + 363. ISBN: 978-1-108-47819-9, £59.99 (hardback). ISBN: 978-1-108-74709-7. £18.99 (paperback).

— Amanda Rees and Charlotte Sleigh, Human London: Reaktion Books, 2020. Pp. 216. ISBN 978-1-78914-214-3. £12.95 (paperback). British Journal for the History of Science, 54(3):387–389, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/patrick-manning-a-
                                        history-of-humanity-the-
                                        evolution-of-the-human-
                                        system-cambridge-cambridge-
                                        university-press-2020-pp-
                                        xiii-363-isbn-9781108478199-
                                        5999-hardback-isbn-
                                        9781108747097-1899-paperback-
                                        amanda-rees-and-charlotte-sleigh-
                                        human-london-reaktion-books-2020-
                                        pp-216-isbn-9781789142143-1295-
                                        paperback/DF2F4E40EFFBB4B64BDC00C630622640.

Turchetti:2023:RNS

[Tur23] Simone Turchetti. Representing noise: stacked plots and the contrasting diplomatic ambitions of radio astronomy and postpunk. British Journal for the History of Science, 56:225–245, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/representing-noise-
stacked-plots-and-the-
contrasting-diplomatic-
ambitions-of-radio-
astronomy-and-postpunk/
085E6BEDC4FD45EE330E23395E387F3F.

vanDijk:2023:BRL

[vD23] Alexander van Dijk. Book reviews: Lydia Barnett, After the Flood: Imagining the Global Environment in Early Modern Europe. Baltimore: Johns Hopkins University Press, 2019. Pp. 264. ISBN 978-1-4214-2951-9. $52.00 (hardback). ISBN 978-1-4214-4527-4. $28.95 (paperback). British Journal for the History of Science, 56 (1):121–123, March 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
lydia-barnett-after-the-
flood-imagining-the-global-
environment-in-early-modern-
europe-baltimore-johns-
hopkins-university-press-
2019-pp-264-isbn-
9781421429519-5200-hardback-isbn-
9781421445274-2895-paperback/
D12F3BF1FF54941FE34C4A5C935778D3.

Weber:2020:BRJ

[Web20] Thomas P. Weber. Book review:

JoAnne Yates and Craig N. Murphy, Engineering Rules: Global Standard Setting since 1880. Baltimore: Johns Hopkins University Press, 2019. Pp. 440. ISBN 978-1-4214-2889-5. $64.95 (hardcover). British Journal for the History of Science, 53(2):283–285, June 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/joanne-yates-and-
craig-n-murphy-engineering-
rules-global-standard-
setting-since-1880-
baltimore-johns-hopkins-
university-press-2019-pp-
440-isbn-9781421428895-6495-
hardcover/8071585BF0754D4D41D014FE278E0C79.

Weber:2023:BRT

[Web23] Thomas P. Weber. Book review: Trais Pearson, Sovereign Necropolis: The Politics of Death in Semi-colonial Siam/ Ithaca, NY: Cornell University Press, 2020. Pp. 233. ISBN 978-1-5017-4015-

2. $49.95 (hardback). British Journal for the History of Science, 56(1):126–128, March 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
trais-pearson-sovereign-
necropolis-the-politics-of-
death-in-semicolonial-siam-
ithaca-ny-cornell-
university-press-2020-pp-
233-isbn-9781501740152-4995-
hardback/
C8013ABE1A4F11556573656D73DF1D0E.

Wesner:2021:RES

[Wes21] Samantha Wesner. Revolutionary electricity in 1790: shock, consensus, and the birth of a political metaphor. British Journal for the History of Science, 54(3):257–275, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
revolutionary-electricity-
in-1790-shock-consensus-and-
the-birth-of-a-political-
metaphor/
1D8ABB62212B1BB32AA805DD543D8EC0.

Widmalm:2023:SBW

[Wid23] Sven Widmalm. ‘Super Bowl of the world conference circuit’? A network approach to high-level science and policy conferencing. British Journal for the History of Science, 56(4):535–551, December 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/super-bowl-of-the-
world-conference-circuit-a-
network-approach-to-
highlevel-science-and-
policy-conferencing/
9AA1752861077BDED686E99732841139.

Wills:2020:BRB

[Wil20] Hannah Wills. Book review:

Benjamin Wardhaugh, Gunpowder and Geometry: The Life of Charles Hutton, Pit Boy, Mathematician, and Scientific Rebel. London: William Collins, 2019. pp. 312. ISBN 978-0-00-829995-8. £20.00/$39.99 (hardback). British Journal for the History of Science, 53(1):129–131, March 2020. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
benjamin-wardhaugh-
gunpowder-and-geometry-the-
life-of-charles-hutton-pit-
boy-mathematician-and-
scientific-rebel-london-
william-collins-2019-pp-312-
isbn-9780008299958-20003999-
hardback/DE0F963903D0DE148EBE4495D6B11E3E.

Williamson:2021:FAA

[Wil21a] Fiona Williamson. Framing Asian atmospheres: imperial weather science and the problem of the local c. 1880-1950. British Journal for the History of Science, 54(3):301–304, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
framing-asian-atmospheres-
imperial-weather-science-
and-the-problem-of-the-
local-c-18801950/
7BED570465846FF2ABB5FE578F414578.

Williamson:2021:JDT

[Wil21b] Fiona Williamson. Just doing their job: the hidden meteorologists of colonial Hong Kong c. 1883–1914. British Journal for the History of Science, 54(3):341– 359, September 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

https://www.cambridge.org/
core/journals/british-
journal-for-the-history-of-
science/article/just-doing-
their-job-the-hidden-
meteorologists-of-colonial-
hong-kong-c18831914/
B3FEAB52A790A800459FC8957A737B08.

Winters:2023:BRJ

[Win23a] Lisa Winters. Book review: J.

David Archibald, Critical Lives: Charles Darwin. London: Reaktion Books, 2021. Pp. 240. ISBN 978-1-78914-440-6. $19.00 (paperback). British Journal for the History of Science, 56(3):420–421, September 2023. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https://www.

cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
j-david-archibald-critical-
lives-charles-darwin-london-
reaktion-books-2021-pp-240-
isbn-9781789144406-1900-
paperback/
E5B3E6744516E02E46CA5EFBE30FE462.

Winters:2023:BRP

[Win23b] Lisa Winters. Book review: Peter H. Hoffenberg, A Science of Our Own: Exhibitions and the Rise of Australian Public Science. Pittsburgh: University of Pittsburgh Press, 2019. Pp. 206.

ISBN 978-0-8229-4576-5. $45.00 (cloth). British Journal for the History of Science, 56:276–277, June 2023. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

                                        /www.cambridge.org/core/
                                        journals/british-journal-
                                        for-the-history-of-science/
                                        article/peter-h-hoffenberg-
                                        a-science-of-our-own-
                                        exhibitions-and-the-rise-of-
                                        australian-public-science-
                                        pittsburgh-university-of-
                                        pittsburgh-press-2019-pp-
                                        206-isbn-9780822945765-4500-
                                        cloth/1FF3BFC7649127CA8B9FE6A904D51F65.

Worboys:2022:IEB

[Wor22] Michael Worboys. Imperial entomology: Boris P. Uvarov and locusts, c. 1920–c. 1950. British Journal for the History of Science, 55(1):27–51, March 2022. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/british-
                                        journal-for-the-history-of-
                                        science/article/imperial-
                                        entomology-boris-p-uvarov-
                                        and-locusts-c1920c1950/
                                        0B29C73859E2E6490020ED0B26956827.

Wynter:2021:AFL

[Wyn21] Rebecca Wynter. Ambition, ‘failure’ and the laboratory: Birmingham as a centre of twentiethcentury British scientific psychiatry. British Journal for the History of Science, 54(1):19–40, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/ambition-failure-
and-the-laboratory-
birmingham-as-a-centre-of-
twentiethcentury-british-
scientific-psychiatry/
4263E8AE6733A2CAEAB6B72D73A2713B.

Yalcinkaya:2022:GSR

[Yal22] M. Alper Yalçinkaya. Globalizing ‘science and religion’: examples from the late Ottoman Empire. British Journal for the History of Science, 55(4):445–458, December 2022. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:/

/www.cambridge.org/core/
journals/british-journal-
for-the-history-of-science/
article/globalizing-science-
and-religion-examples-from-
the-late-ottoman-empire/
5BDB1E43A0A740A982C772EA5109787F.

Yeon-Jung:2021:UGI

[YJ21]  Ji  Yeon-Jung.

Unfriendly guardians: India’s first nuclear leadership change in 1966. British Journal for the History of Science, 54(1):1–17, March 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL https://www.
cambridge.org/core/journals/
british-journal-for-the-
history-of-science/article/
unfriendly-guardians-indias-
first-nuclear-leadership-
change-in-1966/
AAB306F55A8383C4C70B50C19F754A97.

Zaidi:2023:PSC

[Zai23] Waqar H. Zaidi. The Pugwash scientists’ conferences, Cyrus Eaton and the clash of internationalisms, 1954–1961. British Journal for the History of Science, 56(4): 503–517, December 2023. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/journals/
                                        british-journal-for-the-
                                        history-of-science/article/
                                        pugwash-scientists-
                                        conferences-cyrus-eaton-and-
                                        the-clash-of-
                                        internationalisms-19541961/
                                        17262B4D46E0F113F4B97A940294C654.